### v1

# An Exploratory Study on Using Large Language Models for Mutation Testing

Abstract.
The question of how to generate high-utility mutations, to be used for testing purposes, forms a key challenge in mutation testing literature. Large Language Models (LLMs) have shown great potential in code-related tasks but their utility in mutation testing remains unexplored. To this end, we systematically investigate the performance of LLMs in generating effective mutations w.r.t. to their usability, fault detection potential, and relationship with real bugs. In particular, we perform a large-scale empirical study involving 4 LLMs, including both open- and closed-source models, and 440 real bugs on two Java benchmarks. We find that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs, which leads to approximately 18% higher fault detection than current approaches (i.e., 87% vs. 69%) in a newly collected set of bugs, purposely selected for evaluating learning-based approaches, i.e., mitigating potential data leakage concerns. Additionally, we explore alternative prompt engineering strategies and the root causes of uncompilable mutations, produced by the LLMs, and provide valuable insights for the use of LLMs in the context of mutation testing.
Mutation Testing, Mutation Generation, Large Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software testing and debugging
1. Introduction
Mutation testing is one of the most effective software testing techniques (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). It is performed by producing program variants by making simple syntactic changes (i.e., mutations) in the source code of the origin program (i.e., mutants). These variants form the objectives of testing, in a sense requesting tests to trigger different behaviors on the original and the mutant programs. A mutant is called killed or live depending on whether the test execution leads to a different program output (observable behavior) from that of the original program or not. The proportion of the killed mutants over the entire set of mutants forms a test adequacy metric that is called mutation score.
Since mutations are the objectives of testing, the effectiveness of the method strongly depends on the set of mutants that one is using. Traditional techniques use simple syntactic transformations (rules), called mutation operators, that introduce syntactic changes in every code location that they apply (Jia and Harman, 2010; Papadakis et al., 2019). We call these approaches as rule-based approaches.
Unfortunately, rule-based approaches produce large numbers of redundant mutants, i.e., mutants that do not contribute to the testing process (Papadakis et al., 2016; Ammann et al., 2014) and thus exacerbate the method’s computational overheads as well as the human effort involved in examining live mutants. To address this issue, researchers leverage deep learning approaches to form mutation operators based on the project’s history (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022), with the hope that these will produce few but effective mutations. While interesting, these approaches fail to produce effective mutations (Ojdanic et al., 2023b), perhaps due to the fact that they are trained on a few data instances.
Nevertheless, the question of which mutants should be used when performing mutation testing remains open due to the following challenges: (1) Forming effective mutations. To be effective, mutations need to be syntactically correct, which is not the case of most of the existing learning-based methods, and be likely fault revealing (Chekam et al., 2020; Wen et al., 2019), i.e., have a good potential to couple with real faults. While existing methods are effective, there is still significant room for improvement (Degiovanni and Papadakis, 2022). (2) Forming natural mutations. Mutations need to align with natural coding patterns and practices so that they align with developer expectations, i.e., making them feel that they are worth being addressed (Beller et al., 2021). It is a challenge to ensure that the mutated code is not only syntactically correct but also aligns well with the realistic, common coding patterns and practices observed among developers. (3) Forming diverse and killable mutations. Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and to cover the entire spectrum of code and program behaviours.
To deal with these challenges, we aim at using Large-Language Models (LLMs), which have been trained on big code and can produce human-like code (Chen et al., 2021). Following existing studies (Guo et al., 2024; Ma et al., 2024; Schäfer et al., 2023), we designed a default prompt that contains task instructions, the code element to be mutated, the corresponding method code, and a few-shot examples sampled from real-world bugs. We carried out a large-scale evaluation involving 4 LLMs, 2 closed-source commercial and 2 open-source ones, 440 real-world bugs from two Java benchmarks, i.e., Defects4J (Just et al., 2014a) and ConDefects (Wu et al., 2023b), and compared with existing techniques and tools.
Our evaluation includes both quantitative and qualitative analysis for the mutations generated by LLMs, comparing them with existing approaches to assess their cost and effectiveness. Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4-Turbo, CodeLlama-13b-Instruct (Roziere et al., 2023), and StarChat-$\beta$-16b (Li et al., 2023a)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
Our results reveal that the best model for mutant generation is GPT-4 as it outperforms the other LLMs in all metrics used, i.e., produces less equivalent mutants, mutants with higher fault detection potential, as well as higher coupling and semantic similarity with real faults. GPT-3.5 is the second best in terms of fault detection potential, leading to a detection of 96.7% of bugs in Defects4J and 86.7% in ConDefects. From the rest of the techniques, Major is the best, with the potential to reveal 91.6% of bugs in Defects4J and 68.9% in ConDefects, indicating an advantage of approximately 18% for the GPT on ConDefects which is a newly and unseen dataset.
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPTs exhibit the greatest diversity, newly introducing 45 different AST node types, in contrast to just 2 used by traditional approaches such as Major, and subsuming all existing approaches, indicating a much better contextualization of the LLM-based approach.
We also analyzed the mutations generated by GPT that fail to compile, a common issue with learning-based mutations, and found that they fall in 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. This finding indicates that future approaches could focus on directing the LLMs in choosing the right method invocations when mutating.
We wrapped up our method into a mutation tool called Kumo, an LLM-based mutation generation tool for Java. Based on the experimental data, we have also built a dataset that contains high-quality LLM-generated mutants that can be used not only in mutation testing but also in other bug-seeding applications, such as fault localization and fault prediction. The implementation of Kumo and the experimental data are available for open science (kum, 2024).
In summary, our paper makes the following main contributions:
  • We investigate the applicability of LLMs in mutation testing. We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that GPT models excel in generating mutations that closely mimic real bugs.
  • We compare different prompts, and find that few-shot learning with suitable code context archives best performance.
  • We analyze the error types of non-compilable mutations and figure out that method invocation and member assessment are more likely to lead LLMs to generate non-compilable mutations.
  • We build a high-quality dataset of Java mutations that includes manually analysed equivalent mutants.
2. Background and Related Work
Mutation Testing and Mutation Generation.  Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Gligoric et al., 2013; Linares-Vásquez et al., 2017; Wang et al., 2022; Qian et al., 2022; Kazerouni et al., 2021; Groce et al., 2021; Wu et al., 2023a). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011; Zhu et al., 2023; Jiang et al., 2018), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one popular mutation operator is to change one arithmetic binary operator with another one (e.g., $a+b\mapsto a-b$). The majority of existing mutation testing approaches are rule-based, such as PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanch (Schuler and Zeller, 2009), Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), AccMut (Wang et al., 2017), WinMut (Wang et al., 2021), etc. Then researchers proposed using deep learning techniques to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed $\mu$Bert (Degiovanni and Papadakis, 2022). $\mu$Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation (Tip et al., 2024; DEB et al., 2024). However, they have not been thoroughly evaluated, as they only count the number of mutations they introduce and compute the mutation scores achieved by some test suites. In contrast, we thoroughly assess the utility of the mutations we are using (through both a quantitative and qualitative approach), contrast it with the key previous baseline methods, and address all experimental parameters and concerns related to the use of LLMs, i.e., data leakage, cost, and added value.
LLMs for Code.  Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo and GPT-4-Turbo are closed-source LLMs released by OpenAI (Achiam et al., 2023), while CodeLlama and StarChat are open-source. An open-source LLM is usually fine-tuned against a general LLM for specified tasks. For example, CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets, while StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
3. Study Design
3.1. Overview and Research Questions
Our paper intends to thoroughly investigate the capability of existing LLMs in mutation generation. We design research questions covering the evaluation of the default setting of our study, the impact of prompt engineering strategies and models, and root cause analysis for underperforming aspects. Specifically, we aim to answer the following five research questions.
  • RQ1: How do the LLMs perform in generating mutations in terms of cost and usability?
  • RQ2: How do the LLMs perform in generating mutations in terms of behavior similarity with real bugs?
  • RQ3: How do different prompts affect the performance in generating mutations?
  • RQ4: How do different LLMs affect the performance in generating mutations?
  • RQ5: What are the root causes and compilation error types for the generation of non-compilable mutations?
3.2. Datasets
Table 1. Real Bugs Used in Our Experiment
Dataset Project # of Bugs Time Span
Defects4J Math 106 2006/06/05 - 2013/08/31
Lang 65 2006/07/16 - 2013/07/07
Chart 26 2007/07/06 - 2010/02/09
Time 27 2010/10/27 - 2013/12/02
Closure 133 2009/11/12 - 2013/10/23
Mockito 38 2009/06/20 - 2015/05/20
ConDefects 45 2023/09/01 - 2023/09/30
Total 440 2006/07/16 - 2023/09/30
We intend to evaluate our approach with real bugs, and thus we need to use bug datasets with the following properties:
  • The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
  • The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
  • Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
  • Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J v1.2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2023b) to evaluate the mutation generation approaches, shown as Table 1. In total, we conduct the experiments on 440 bugs.
Defects4J is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023a; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from 6 open-source projects of diverse domains, ensuring a broad representation of real-world bugs. In total, these 6 projects contain 395 real bugs. However, from Table 2 and Table 1, we observe the time spans of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2023b), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder (atc, 2024) programming contest. To prevent data leakage, we exclusively use bugs reported after the LLMs’ release date, specifically those identified on or after August 31, 2023, and in total we collect 45 Java programs from ConDefects.
3.3. Mutation Generation via LLMs
Given the location of a Java program, we extract essential information (e.g., the context method or the corresponding unit tests) to formulate prompts instructing mutation generation, then feed prompts to the selected LLM. Once the LLM returns its responses, we filter out mutations that are non-compilable, redundant, or identical to the original code, and return the generated mutations. As shown in Section 3.4, our study supports a comprehensive set of metrics for evaluating mutations.
3.3.1. Models
We aim to comprehensively compare the existing LLMs in their ability to generate mutations from various perspectives. This includes evaluating the capabilities of models that are fine-tuned on different base models and contrasting the performance of commercial closed-source models with open-source alternatives. In our study, we select LLMs based on the following criteria. (1) The LLM must have coding capabilities, i.e., understanding, generating, and transforming code. (2) The LLM must support interaction in chat mode, allowing for conversational code transformation. (3) The LLM must offer accessible APIs for integration.
Finally, we select the following models: GPT-3.5-Turbo (Achiam et al., 2023), GPT-4-Turbo (Achiam et al., 2023), CodeLlama-13b-Instruct (Roziere et al., 2023), and StarChat-$\beta$-16b (Li et al., 2023a)). The details of these models are shown as Table 2. These LLMs cover open-source (i.e., StarChat and CodeLlama) and closed-source (i.e., GPT-3.5-Turbo and GPT-4-Turbo), and various kinds of base models.
Table 2. Studied LLMs
Model Type Model Training Data Time Release Time Size
Closed GPT-3.5-turbo 2021/09 2023/03 20B
GPT-4-turbo 2023/04 2023/06 110B
Open CodeLlama-13b-Instruct 2023/08 13B
StarChat-$\beta$-16b 2023/06 16B
3.3.2. Prompts
Prompts are crucial for LLMs because they guide the models’ responses and determine the direction of the generated output. A well-crafted prompt can significantly influence the utility, relevance, and accuracy of the code produced by an LLM. This subsection introduces how we design the prompts.
To design effective prompts for mutation generation, we follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experimentation. Figure 1 shows the default prompt template in our study.
In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2021, 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Following the existing study on few-shot examples (Deng et al., 2024), we sample 6 of the bugs from QuixBugs.
To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all 6 examples. Table 3 shows these examples.
Table 3. Default Few-Shot Examples from QuixBugs
Correct Version Buggy Version
n = (n & (n - 1)); n = (n ^ (n - 1));
while (!queue.isEmpty()) while (true)
return depth==0; return true;
ArrayList r = new ArrayList();
r.add(first).addll(subset);
to_add(r);
to_add.addAll(subset);
c = bin_op.apply(b,a); c = bin_op.apply(a,b);
while(Math.abs(x-approx*approx)>epsilon) while(Math.abs(x-approx)>epsilon)
Figure 1. The Default Prompt Template
3.4. Evaluation Metrics
To assess the utility of the LLM-generated mutations, we should define a set of metrics that comprehensively evaluate them. Therefore, we designed several categories of metrics capturing different aspects as follows: Cost Metrics, Usability Metrics, and Behavior Metrics. Besides these metrics, we also additionally report the Mutation Count and Mutation Score. Note that both of them cannot directly reflect the quality of generated mutations. For example, a higher mutation score does not indicate a higher quality, which may also be caused by generating a large proportion of mutants that are easily killed.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set $A$, where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set $C$. Note that the set $C$ contains two types of mutations that should be removed, the useless mutations (denoted as $U$) and equivalent mutations (denoted as $E$). The useless mutations refer to those that are identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set $(C-E-U)$, excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set $E$ is undecidable, so we follow the common practice and use the set $(C-U)$ as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022).
3.4.1. Cost Metrics
The Cost Metrics indicate time and economic costs in mutation generation, which contains the following two metrics.
Average Generation Time:  the average duration needed to produce each mutation, measuring the efficiency of an approach.
Cost per 1K Mutations:  the average money cost in US dollars to generate 1k mutations. For GPT-3.5 and GPT-4, we calculate the average cost by API usage, while for the open-source models, we employ the costs of renting cloud servers.
3.4.2. Usability Metrics
The Usability Metrics serve as key indicators of the practical utility of generated mutations.
Compilability Rate:  the proportion of mutations that successfully compile, computed as: ${|C|}/{|A|}$. Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
Useless Mutation Rate:  the proportion of useless mutations, computed as: $|U|/|A|$. LLMs sometimes generate duplicate mutations or mutations that are exactly the same as the original code. While traditional approaches rarely generate useless mutations, LLMs suffer from issues of misunderstanding and repetition, leading to generating these useless mutations.
Equivalent Mutation Rate:  the proportion of equivalent mutation, computed as: $|E|/|A|$. Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. Due to the undecidable nature of identifying whether a mutation is equivalent, we sample significant portions of all mutations and judge them manually. More specifically, given the population to be the set $A$, we set the confidence level to be 95% and the margin of error to be 5%, to get a subset of mutations to be manually inspected.
AST Node Diversity:  We prefer the set of mutations with diverse forms, therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior.
3.4.3. Behavior Similarity with Real Bugs
Behavior similarity refers to the degree of similarity in functionality and execution between a mutation and its corresponding real bug. Because capturing all program behaviors is undecidable, following existing studies, we thus employ the following metrics as a proxy to measure behavior similarity (Ojdanic et al., 2023a; Papadakis et al., 2018; Khanfir et al., 2023).
Real Bug Detectability:  Given the location of a real bug (which may occupy multiple lines), let a mutation generation approach generate a set of mutations, if the tests that kill these mutations can also kill this real bug, we say that the mutation generation approach can “detect” this real bug. Given a set of real bugs and a mutation generation approach, Real Bug Detectability quantifies the number of bugs detected by this mutation approach.
Coupling Rate:  Mutation testing is based on the fault coupling effect, i.e., mutations though are simple syntactic changes, coupled to more complex bugs (Offutt, 1992). To describe the extent to which a mutation couples to its corresponding real bug, we adopt the Coupling Rate following the existing study (Just et al., 2014b). Given a real bug and a corresponding mutant generated at the same location, the mutant is deemed a coupled mutant if the set of the test cases killing the mutant overlaps with the set of the bug-triggering test cases. The Coupling Rate refers to the proportion of coupled mutants within $(C-U)$.
Ochiai Coefficient:  approximates the semantic similarity by passing and failing test cases (Offutt and Untch, 2001), which is also widely used in the field of fault localization (Zou et al., 2019). Following the existing study (Ojdanic et al., 2023a), we employ the Ochiai Coefficient to calculate how a mutation resembles its corresponding real bug. More concretely, let $M$ and $B$ be a mutation and its corresponding real bug, and let $fT_{M}$ and $fT_{B}$ be the set of their failing test cases respectively. $M$’ coefficient is computed by the following formula:
(1) $Ochiai(M,B)=\frac{|fT_{M}\cap fT_{B}|}{\sqrt{|fT_{M}|\times|fT_{B}|}}$
Therefore, if $M$’s Ochiai Coefficient is $1$, it is semantically equal to its corresponding bug $B$.
3.5. Experiment Settings
3.5.1. Settings for Mutation Generation
One major target of our study is to investigate the similarity between LLM-generated mutations and real bugs. Therefore, we generate mutations within the context of real bugs to enable such a comparison. To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the full set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 5.
3.5.2. Settings for RQ1-RQ2
In the first part of our study, we intend to qualitatively and comparatively evaluate the capability of LLMs in generating mutations. We adopt the default settings, which contain two models (i.e., GPT-3.5-Turbo and CodeLlama), and use the default prompt template, generating mutations on the universal set of the 405 bugs in Section 3.2.
To understand the performance of the LLMs, we adopt LEAM (Tian et al., 2022), $\mu$Bert (Degiovanni and Papadakis, 2022), PIT (Coles et al., 2016), and Major (Just et al., 2011) as baselines. Among these approaches, LEAM is the state-of-the-art mutation generation approach which employs a model trained from 13 million real Java bugs. While $\mu$Bert is based on Bert (Devlin et al., 2018), a masked pre-trained language model that can not conversational respond to human-like text and code.PIT (Coles et al., 2016) and Major (Just et al., 2011) are the popular traditional mutation testing tools with human-predefined simple mutation operators. Note that for PIT and Major, we employ all mutation operators.
3.5.3. Settings for RQ3-RQ4
In the second part of our study, we intend to explore how different prompts and different models impact the performance of mutation generation. Due to budget limitations, we sample a subset of 105 bugs from our dataset, which consists of 10 sampled bugs from each Defects4J project and all ConDefects bugs.
In RQ3, for the prompts exploration, we use the GPT-3.5-Turbo model and switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get 4 types of prompts, shown as follows:
  1. (1)
    Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
  2. (2)
    Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the examples.
  3. (3)
    Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
  4. (4)
    Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
In RQ4, for the LLMs exploration, we use the default prompt and investigate the performance of GPT-3.5-Turbo, GPT-4-Turbo, CodeLlama-13b-Instruct, and StarChat-$\beta$-16b.
3.5.4. Settings for RQ5
Generating non-compilable mutations incurs additional meaningless costs, so in the third part of our study, we study the root causes behind LLM generating non-compilable mutations. Specifically, we sample 384 non-compilable mutations of each mutation generation approach and analyze their reason rejected by the Java compiler and the features of their surrounding code context.
4. Evaluation Results
We performed our experimental analysis on two cloud servers, each of which is equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
Table 4. Overall Performance of All the Mutation Generation Techniques
Metric Category Metric GPT 3.5 CodeLlama-13b LEAM $\mu$BERT PIT Major
D4J CD D4J CD D4J CD D4J CD D4J CD D4J CD
Mut. Score 0.731 0.734 0.716 0.698 0.529 0.682
Mut. Count 351,557 303,166 671,737 141,031 1,032,392 314,694
Cost Avg. Gen. Time 1.79 9.06 3.06 2.23 0.017 0.083
\$ Cost per 1K Mut. 0.288 0.577 0.195
Usability Compilability Rate 61.7% 79.6% 75.8% 73.3% 33.8% 50.9% 21.9% 26.6% 98.3% 92.8%
Useless Mut. Ratio 11.6% 9.2% 41.0% 9.3% 0.5% 9.6% 1.8% 4.6% 0.0% 0.0%
Eq. Mut. Rate 2.3% 2.1% 1.3% 0.7% 1.6% 1.0% 2.1% 3.1% 0.5% 0.8%
Behavior Real Bug Detectability 0.967 0.867 0.906 0.667 0.767 0.489 0.785 0.378 0.511 0.489 0.916 0.689
Coupling Rate 0.416 0.625 0.398 0.612 0.267 0.318 0.409 0.581 0.133 0.593 0.342 0.479
Ochiai Coefficient 0.638 0.689 0.390 0.378 0.286 0.400 0.230 0.578 0.271 0.489 0.519 0.600
4.1. RQ1: Performance on Cost and Usability
4.1.1. Cost
For each technique, we record the generation duration and the cost for API usage or rending GPU servers, shown as the Cost lines of Table 4. In addition, we also report the mutation count and mutation score, although they do not fully reflect the overall quality of the mutation set.
From the table, we can find that GPT-3.5 and CodeLlama-30b-Instruct generate 351,557 and 303,166 mutations111 Although GPT-3.5 and CodeLlama-13b are instructed to generate the same number of mutations, they do not consistently follow the instructions, thereby yielding different numbers of mutations., with a mutation score of 0.731 and 0.734, respectively. Traditional approaches PIT and Major yield the lowest mutation score among all the approaches.
In terms of cost, GPT-3.5 and CodeLlama-13b cost 1.79s and 9.06s to generate mutations, respectively. While LEAM, $\mu$Bert, PIT, and Major cost 3.06s, 2.23s, 0.017s, and 0.083s, respectively, indicating that previous approaches are faster than LLM-based approaches in generating mutations. In terms of economic cost, we only record the API usage cost of GPT-3.5 and the cloud server cost for running CodeLlama and LEAM, because the remaining approaches barely need GPU resources. The result indicates that renting cloud servers to run the open-source models costs more than using API for GPT-3.5.
Finding 1: Previous approaches cost less than LLM-based approaches in generating mutations. For example, in terms of mutation generation time, GPT-3.5 takes 1.79s to generate one mutation, CodeLlama-13b takes 9.06s, while PIT and Major cost 0.017s and 0.083s, respectively. Regarding economy cost, traditional approaches barely need GPU resources, but GPT-3.5 costs 0.288\$ to generate 1,000 mutations using API, and running CodeLlama-13b on a cloud server costs 0.577\$.
4.1.2. Mutation Usability
The usability of mutations (i.e., whether the mutations can be used to calculate mutation score) is affected by the capabilities of LLMs in correctly understanding the task and generating compilable, non-duplicate, and non-equivalent mutations. The results are shown as the Usability rows in Table 4. Due to the potential threat of data leakage, we represent the data of Defects4J and Condefects separately. Note that the mutations of PIT are in the form of Java bytecode and cannot be transformed back to the source level, so we skip PIT when answering this RQ.
Compilability Rate: We observe that GPT-3.5 achieves a compilability rate of 61.7% for Defects4J and 79.6% for ConDefects datasets. While CodeLlama-13b achieves a higher compilability rate of 75.8% for Defects4J and 73.3% for ConDefects. The rule-based approach Major achieves the highest compilabity rate (i.e., 98.3% on Defects4J and 92.8% on ConDefects). Note that though Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true)
Useless Mutation Rate: We also observe that LLMs generate a significant portion of duplicate and useless mutations. For example, for Defects4J, GPT-3.5 generates 11.6% useless mutations, accounting for 18.8% of all compilable mutations (i.e., 11.6%/61.7%). In contrast, LEAM, $\mu$Bert, and Major seldom generate duplicate mutations, which is likely due to these approaches being governed by strict grammar rules that limit the production of redundant mutations. For example, Major iteratively applies different mutation operators producing no redundant mutants.
Table 5. Sample Results for Equivalent Mutations
Dataset Type GPT-3.5 CL-13b LEAM $\mu$Bert Major
Defecs4J # All 351,332 302,941 671,342 140,753 314,527
# Sampled 384 384 384 384 384
# Eq. Mut. 9 5 6 8 2
ConDefects # All 225 225 395 278 167
# Sampled 143 143 196 162 123
# Eq. Mut. 3 1 2 5 1
Equivalent Mutation Rate: Equivalent mutations change the source code syntactically without affecting the program’s functionality, thereby affecting the accuracy of mutation score calculations. Due to the undecidable nature of identifying equivalent mutations, it is not feasible to design an algorithm that can completely detect them. Therefore, we sample a subset of mutations for each technique and manually assess whether they are equivalent. To maintain a 95% confidence level and a 5% margin of error, we sampled a fixed number of mutations from each approach’s output, to calculate the Equivalent Mutation Rate. Following the protocol of existing studies (Islam et al., 2019; Guo et al., 2024), two authors, each with over five years of Java programming experience, first received training on the equivalent mutants, and then independently labeled the mutations. The final Cohen’s Kappa coefficient score (Viera et al., 2005) is more than 90%, indicating a high level of agreement. Table 5 shows the number of sampled mutations and the number of identified equivalent mutations.
We again observe that Major performs the best in generating non-equivalent mutants. In particular, as shown in the bottom row of Usability section in Table 4, for GPT-3.5, 2.3% of mutations generated on Defects4J are equivalent, while the rate is 2.1% on ConDefects. CodeLlama-13b shows lower rates, with 1.3% and 0.7% on Defects4J and ConDefects, respectively. LEAM generates equivalent mutations at rates of 1.6% on Defects4J and 1.0% on ConDefects. $\mu$Bert produces 2.1% on Defects4J and 3.1% on ConDefects. Major achieves the lowest rates, with only 0.5% on Defects4J and 0.8% on ConDefects.
When comparing the two datasets Defects4J and ConDefects for the three usability aspects introduced above, we observe that the metrics vary between the Defects4J and ConDefects datasets for each mutation approach. This variation may be attributed to differences in the complexity and structure of bugs present in each dataset. Mutations in Defects4J may involve APIs and global variables that are not explicitly mentioned in prompts, making the context more complex. This complexity results in both metrics being inferior on Defects4J compared to ConDefects.
Finding 2: Major outperforms LLM-based and other learning-based methods in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, GPT-3.5 and CodeLlama-13b exhibit compilability rates of 61.67% and 75.86%, useless mutation rates of 11.60% and 40.96%, and 2.28% and 1.14% equivalent mutations, respectively. In contrast, Major excels with a 98.33% compilability rate, 0% useless mutations, and 0.59% equivalent mutants.
Figure 2. Number of types and examples of newly introduced AST Nodes from different mutation approaches.
4.1.3. AST Node Dirversity
To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by Javalang (jav, 2024), to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 shows the results, which reveal that LLM-based mutation approaches introduce more types of new AST nodes than other approaches. GPT-3.5 exhibits the greatest diversity, introducing 45 new AST node types, and subsuming all other approaches. Traditional approach Major introduces only two new types of AST nodes, which is not surprising as the mutation operators are a fixed set of simple operations.
We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Few-shot Example” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node. This indicates that the LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
Table 6. Ratio of Deletion Mutations and the Top-3 Newly Introduced AST Node Types for Each Approach.
Top K GPT-3.5 CodeLlama-13b LEAM $\mu$BERT Major
Del. 5.1% 0.3% 49.7% 0.2% 17.6%
1st LT (47.6%) LT (49.5%) LT (23.6%) LT (53.8%) ST (59.9%)
2nd BO (24.7%) BO (35.5%) MI (17.6%) MR (22.7%) LT (40.1%)
3rd MI (8.2%) MI (4.1%) MR (16.3%) MI (12.5%)      —
ST: Statement, BO: BinaryOperation, LT: Literal, MI: MethodInvocation, MR: MemberReference
We further analyzed the distribution of deletions and AST Node types across different approaches, as shown in Table 6, which highlights the top-3 most frequent AST node types. For the proportion of deletions, GPT-3.5 and CodeLlama-13b generate 5.1% and 0.3% deletions, respectively. Nearly half of the mutations of LEAM are deletions, which may be the reason for generating many non-compilable mutations. Major generates deletions on restricted locations and results in 17.6% deletions. $\mu$Bert replaces the masked code elements with BERT, thereby rarely having deletions.
Regarding AST Node types, the most frequent new AST node type of all the learning-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the rule-based approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the Deletion operator, there exists a significant portion of producing empty statements, leading to a large proportion.
Finding 3: LLM-based mutation approaches introduce more types of new AST nodes than other approaches. The mutations generated by GPT-3.5 introduce 45 new types of AST nodes, demonstrating the highest level of diversity.
4.2. RQ2: Behavior Similarity
The bottom three rows of Table 4 present a comparative evaluation of Behavior Metrics for mutation generation approaches.
4.2.1. Real Bug Detectability
GPT-3.5 detects 382 bugs of the 395 Defects4J bugs and 39 bugs of all 45 ConDefects bugs, i.e., 96.7% Defects4J bugs and 86.7% ConDefects bugs could be revealed by these mutations, achieving the best performance. CodeLlama-13b detects 358 bugs from Defects4J (i.e., 90.6%) and 30 bugs from ConDefects (i.e., 66.7%), respectively. Major achieves the second-best performance by detecting 362 Defects4J bugs (i.e., accounting for 91.6%) and 31 ConDefects bugs (i.e., accounting for 68.9%).
Finding 4: GPT-3.5 has the best performance in Real Bug Detectability, identifying 96.7% of bugs in Defects4J and 86.7% in ConDefects. Among the other approaches, Major performs best, detecting 91.6% of bugs in Defects4J and 68.9% in ConDefects.
4.2.2. Coupling Rate
The Coupling Rate measures the degree of coupling between the generated mutations and their corresponding real bugs. GPT-3.5 exhibits coupling rates of 0.416 on Defects4J and 0.625 on ConDefects, respectively, achieving the best performance on both datasets. While CodeLlama-13b achieves coupling rates of 0.398 and 0.612, respectively.
Finding 5: GPT-3.5 leads in Coupling Rate with scores of 0.416 on Defects4J and 0.625 on ConDefects. $\mu$Bert follows closely with 0.409 on Defects4J and 0.581 on ConDefects.
4.2.3. Ochiai Coefficient
The Ochiai Coefficient measures the semantic similarity between mutations and real bugs. GPT-3.5 leads with coefficients of 0.638 on Defects4J and 0.689 on ConDefects, outperforming CodeLlama-13b, which scores 0.39 and 0.378 on the respective datasets. Despite the notable performance gap between the two, their results are consistent across datasets. Major ranks second with coefficients of 0.519 on Defects4J and 0.6 on ConDefects.
Finding 6: GPT-3.5 leads in Ochiai Coefficient performance with scores of 0.519 on Defects4J and 0.6 on ConDefects. Major closely follows, matching these scores and securing the second-best ranking among the approaches.
4.3. RQ3: Impacts of Different Prompts
Table 7. Comparison Results of Different Prompts and LLMs
Metric Category Metric GPT-3.5 Closed Source LLM Open Source LLM
P1 P2 P3 P4 GPT-3.5 GPT-4 StarChat-16b CodeLlama-13b
Mutation Count 47,666 47,905 41,548 39,638 47,666 23,526 25,496 41,808
Mutation Score 0.705 0.703 0.731 0.782 0.705 0.702 0.538 0.734
Cost \$ Cost per 1K Mutations 0.288 0.288 0.287 0.289 0.288 18.701 0.876 0.577
Average Generation Time 1.79 1.78 1.81 1.88 1.79 42.25 13.76 9.06
Usability Compilability Rate 65.7% 64.1% 61.9% 60.8% 65.7% 79.9% 58.2% 74.1%
Useless Mutation Ratio 13.7% 11.3% 22.3% 15.0% 13.7% 7.2% 38.3% 39.2%
Equivalent Mutation Rate 1.5% 1.2% 1.2% 1.5% 1.5% 0.9% 2.1% 1.0%
Behavior Real Bug Detectability 0.895 0.857 0.714 0.629 0.895 0.905 0.324 0.724
Coupling Rate 0.140 0.136 0.096 0.096 0.140 0.143 0.101 0.110
Ochiai Coefficient 0.590 0.571 0.438 0.362 0.590 0.619 0.152 0.286
The left half of Table 7 presents the comparative results of GPT-3.5 via different prompts listed in Section 3.5.3. Prompts P1 to P3 progressively simplify, each containing less information than its predecessor, while P4 is the most complex, enhancing P1 with test suite codes.
Overall, P1, the default prompt, excels in Compilability Rate and all Behavior Metrics. P2, created by removing few-shot examples from P1, leads in Average Generation Time, Useless Mutation Ratio, and Equivalent Mutation Rate, suggesting improved quality in compilable mutations. P3, provided only with the code element to be mutated, achieves the lowest cost for using the least tokens. Conversely, P4, which extends P1 with test suite codes, shows the lowest performance across all metrics, suggesting that GPT-3.5 cannot effectively utilize test suite data to enhance mutation quality.
Finding 7: The prompt with the whole method and few-shot examples as context (i.e., P1) archives the best performance across all behavior metrics, whereas adding the code of test suites (i.e., P4) decreases the performance.
4.4. RQ4: Impacts of Different LLMs
To answer this RQ, we add two extra LLMs, GPT-4 and StarChat-16b, and compare their results with the two default LLMs, GPT-3.5 and Code Llama-13b. The right half of Table 7 shows comparative results across models using the default prompt. We observe that closed-source LLMs generally outperform others in most metrics. GPT-3.5 excels in Mutation Count, Generation Cost per 1K Mutations, and Average Generation Time, ideal for rapidly generating numerous mutations. GPT-4 leads in all Usability Metrics, and Behavior Similarity Metrics, demonstrating its effectiveness in code-related tasks, although its enhancements over GPT-3.5 in Behavior Metrics are trivial. Between the two open-source LLMs, despite StarChat-$\beta$-16b having more parameters, CodeLlama-13b outperforms in all metrics. This suggests that model architecture and training data quality significantly impact performance beyond just the number of parameters.
Finding 8: Among the four LLMs we study, GPT-3.5 excels in generating a large number of mutations with lower costs, while GPT-4 outperforms in all the Usability Metrics, and Behavior Similarity Metrics.
4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 4.1, LLMs generate a significant number of non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT-3.5, ensuring the confidence level is 95% and the margin of error is 5%. From the manual analysis of these non-compilable mutations, we identified 9 distinct error types, as shown in Table 8.
Table 8. Error Types of Non-Compilable Mutations
ID Error Type % ID Error Type %
1 Usage of Unknown Methods 27.3% 6 Type Mismatch 7.6%
2 Code Structural Destruction 22.9% 7 Incorrect Initialization 3.1%
3 Incorrect Method Parameters 12.8% 8 Incorrect Location 3.1%
4 Usage of Unknown Variables 11.2% 9 Incorrect Exceptions 2.3%
5 Usage of Unknown Types 9.6%
Shown as Table 8, the most common error type, Usage of Unknown Methods, accounts for 27.34% of the total errors, revealing the hallucination issue of generative models (Huang et al., 2023). Code Structural Destruction is the second most common error, accounting for 22.92%, indicating that ensuring generated codes are syntactically correct remains a challenge for current LLMs. This result suggests that there is still significant room for improvement in current LLMs.
Finding 9: The mutations generated by GPT-3.5 have 9 main compilation error types, with Usage of Unknown Methods and Code Structural Destruction being the top-2 most prevalent types.
(a) GPT-3.5
(b) CodeLlama-13b
(c) LEAM
(d) $\mu$Bert
(e) Major
Figure 3. Distribution of the AST Node Types of Origin Code for Non-Compilable Mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by GPT-3.5, CodeLlama, LEAM, and $\mu$Bert in Section 4.1, as shown in Figure 3. For all approaches, the code locations with MethodInvocation and MemberReference are the top-2 prevent AST Node types. In particular, there are more than 30% of non-compilable mutations occur at the location with MethodInvocation, and 20% occur at the location with MemberReference. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that for GPT-3.5, CodeLlama, LEAM, $\mu$Bert, and Major, these mutations account for 7.1%, 0.2%, 45.3%, 0.14%, and 14.4% of all their non-compilable mutations, respectively. Thus for LLMs, deletion is not the major reason for non-compilation.
Finding 10: Method invocations and member references in the code are the most likely to trigger LLMs to generate non-compilable mutations.
5. Discussion
5.1. Sensitivity to Chosen Experiment Settings
5.1.1. Experiment Setup of the Context Length
Table 9. Performance of Different Context Lengths
Metric Category Metric 1-Line Context 2-Line Context 3-Line Context
Usability Compilability Rate 0.610 0.617 0.657
Useless Mutation Ratio 0.119 0.128 0.137
Equivalent Mutation Rate 1.6% 1.1% 1.3%
Behavior Real Bug Detectability 0.933 0.917 0.917
Coupling Rate 0.141 0.121 0.136
Ochiai Coefficient 0.517 0.417 0.467
Table 10. Performance Similarity of Different Context Lengths
Settings Spea. Coef. $p$-value Pear. Coef. $p$-value
1-Line v.s. 2-Line 0.9762 3.31E-05 0.9980 2.12E-08
2-Line v.s. 3-Line 0.9762 3.31E-05 0.9997 1.04E-10
1-Line v.s. 3-Line 0.9524 2.60E-04 0.9978 2.69E-08
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 9. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in Table 10. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
5.1.2. Impact of Different Few-Shot Examples
Table 11. Impact on Different Few-Shot Examples
Metric QB-6a (Default) QB-3 QB-6b QB-9 D4J-6
Comp. Rate 65.7% 62.7% 63.9% 61.8% 62.1%
Useless Mut. Ratio 13.7% 12.8% 12.8% 13.8% 12.5%
Equivalent Mutation Rate 1.3% 1.6% 1.3% 1.1% 1.6%
Real Bug Det. 0.917 0.933 0.917 0.917 0.900
Coupling Rate 0.136 0.128 0.132 0.121 0.139
Ochiai Coefficient 0.467 0.450 0.434 0.434 0.367
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples.
Table 12. Performance Similarity of Different Few-shot Examples
Settings Spea. Coef. $p$-value Pear. Coef $p$-value
QB-6a (Default) v.s. QB-3 0.9701 6.55E-05 0.9996 1.70E-10
QB-6a (Default) v.s. QB-3b 0.9762 3.31E-05 0.9984 1.00E-08
QB-6a (Default) v.s QB-9 0.9762 3.31E-05 0.9985 7.98E-09
QB-6a (Default) v.s D4J-6 0.9524 2.60E-04 0.9980 2.13E-08
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 11.
We further compute their similarity, shown as Table 12. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude the performance of different prompts is very similar.
5.1.3. Performance of Using the Same Number of Mutations
In Section 4, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 16,785. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 13. Therefore, under fixed quantity conditions, the mutations generated by GPT-3.5-Turbo are still the closest to real bugs in behavior.
Table 13. Performance Under the Same Number of Mutations
Metric GPT-3.5 CL-13b LEAM $\mu$BERT Major PIT
Comp. Rate 62.9% 74.2% 38.9% 18.5% 96.5%
Useless Mut. Ratio 11.0% 39.0% 1.4% 1.4% 0.0%
Real Bug Det. 0.892 0.735 0.675 0.550 0.833 0.437
Coupling Rate 0.133 0.105 0.132 0.132 0.128 0.024
Ochiai Coefficient 0.358 0.178 0.167 0.150 0.341 0.192
5.2. Implications
Based on our findings, we discuss the pros and cons of using LLMs for generating mutations and suggest potential improvements.
First, our study shows that existing LLMs have the potential to generate high-quality mutations for mutation testing with acceptable costs. The greatest advantage of LLMs is their ability to generate mutations that more closely mimic real bugs in behavior. For example, Listing 1 shows two compilable mutations generated by GPT-3.5 in Chart-1, which are theoretically beyond the capabilities of existing approaches. In Mutation-1, the LLM inferences there must exist leftBlock because of the occurrence of rightBlock, and replaces the variable correspondingly. This mutation is not killed by the bug-triggering test. Without natural language understanding capabilities (i.e., understanding the correspondence between “right” and “left”), it is impossible to create such mutations. In Mutation-2, the LLM modifies the else branch to else if, employing an integer variable seriesCount from context to synthesize the condition, showing the capabilities of code structural changing that existing methods cannot achieve. Also, this mutation is alive and can guide test enhancement.
// Mutation-1
-   Size2D size = this.rightBlock.arrange(g2, c4);
+   Size2D size = this.leftBlock.arrange(g2, c4);
// Mutation-2
-   else {
-     g2.setStroke(getItemOutlineStroke(row, column, selected));
+   else if (seriesCount == 1) {
+     g2.setStroke(getItemOutlineStroke(row, column, selected));
Listing 1: Two Example Mutations of GPT-3.5 on Chart-1
Second, our study explores how different prompts, models, contexts, and few-shot examples impact mutation generation performance. Our study highlights the importance of using proper prompts and models with stronger coding capabilities. In particular, designing prompts requires providing essential context information (such as the whole method), but including too much information (such as the target unit test) may reduce effectiveness.
Third, although LLMs achieve promising results in mutation testing, our study indicates that there is still significant room for improvement. Specifically, current LLMs generate a substantial number of mutations that are not compilable. Therefore, we analyze the types of mutations that lead to non-compilable code and identify some root causes of the LLMs. We advocate for more research efforts in program analysis and repair approaches to address error-prone code elements for mutation such as method invocations.
5.3. Threats to Validity
The selected LLMs, programming language, datasets, and baseline approaches could be a threat to the validity of our results. To mitigate this threat, we adopt the most widely studied models (i.e., GPT and CodeLlama), the most popular language (i.e., Java), and the most popular dataset Defects4J. We also employ state-of-the-art mutation testing approaches as baselines, including learning-based (i.e., $\mu$Bert and LEAM) and rule-based (i.e., PIT and Major).
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2023b) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. Our results are: GPT, CodeLlama, Major, LEAM, and $\mu$Bert, 282, 77, 67, 386, 39 on the Defects4J dataset while on the ConDefects are 7, 9, 13, 8, 1, respectively, and indicate that on Defects4J GPT and LEAM, approaches tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert wields significantly the least number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPT and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interesting, in the ConDefects dataset, which has not been seen by any of the tools, Major has the majority of the exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address this threat, we elaborately explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar.
The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
6. Conclusion
In this paper, we systematically investigate the performance of LLMs in generating mutations. We evaluate their utility from several aspects and find LLMs have the advantage of generating diverse mutations that mimic the behaviour of real bugs. We also analyze and discuss some limitations of LLMs and point out further directions. We advocate research efforts to promote LLM for mutation testing in the future.