### v2

# On the Use of Large Language Models in Mutation Testing

Abstract.
Mutation testing is a foundation approach in the software testing field, based on automatically seeded small syntactic changes, known as mutations. The question of how to generate high-utility mutations, to be used for testing purposes, forms a key challenge in mutation testing literature. Large Language Models (LLMs) have shown great potential in code-related tasks but their utility in mutation testing remains unexplored. To this end, we systematically investigate the performance of LLMs in generating effective mutations w.r.t. to their usability, fault detection potential, and relationship with real bugs. In particular, we perform a large-scale empirical study involving six LLMs, including both state-of-the-art open- and closed-source models, and 851 real bugs on two Java benchmarks (i.e., 605 bugs from 12 projects of Defects4J 2.0 and 246 bugs of ConDefects). We find that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs, which leads to approximately 19% higher fault detection than current approaches (i.e., 93% vs. 74%). Nevertheless, the mutants generated by LLMs have worse compilability rate, useless mutation rate, and equivalent mutation rate than those generated by rule-based approaches. This paper also examines alternative prompt engineering strategies and identifies the root causes of uncompilable mutations, providing insights for researchers to further enhance the performance of LLMs in mutation testing.
Mutation Testing, Mutation Generation, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Automated Software Engineering; 27 October - 1 November, 2024; Sacramento, CA, USAisbn: 978-1-4503-XXXX-X/XX/XXjournal: JACMjournalvolume: 1journalnumber: FSEarticle: 1publicationmonth: 6ccs: Software and its engineering Software testing and debugging
1. Introduction
Mutation testing is one of the most widely-studied software testing techniques (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). It is performed by producing program variants by making simple syntactic changes (i.e., mutations) in the source code of the origin program (i.e., mutants). These variants form the objectives of testing, in a sense requesting tests to trigger different behaviors on the original and the mutant programs. A mutant is called killed or live depending on whether the test execution leads to a different program output (observable behavior) from that of the original program or not. The proportion of the killed mutants over the entire set of mutants forms a test adequacy metric that is called mutation score.
Since mutations are the objectives of testing, the effectiveness of the method strongly depends on the set of mutants that one is using. Traditional techniques use simple syntactic transformations (rules), called mutation operators, that introduce syntactic changes in every code location that they apply (Jia and Harman, 2010; Papadakis et al., 2019). We call these approaches as rule-based approaches.
Unfortunately, rule-based approaches produce large numbers of redundant mutants, i.e., mutants that do not contribute to the testing process (Papadakis et al., 2016; Ammann et al., 2014) and thus exacerbate the method’s computational overheads as well as the human effort involved in examining live mutants. To address this issue, researchers leverage deep learning approaches to form mutation operators based on the project’s history (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., [n. d.]), with the hope that these will produce few but effective mutations. While interesting, these approaches still have limitations in producing effective mutations (Ojdanic et al., 2023b), perhaps due to the fact that they are trained on a limited number of data instances.
To date, the question of how to generate powerful mutants remains open due to the following challenges: (1) Forming effective mutations. To be effective, mutations need to be syntactically correct, which is not the case of most of the existing learning-based methods, and be likely fault revealing (Chekam et al., 2020; Wen et al., 2019), i.e., have a good potential to couple with real faults. While existing methods are effective, there is still significant room for improvement (Degiovanni and Papadakis, 2022). (2) Forming natural mutations. Mutations need to align with natural coding patterns and practices so that they align with developer expectations, i.e., making them feel that they are worth being addressed (Beller et al., 2021). It is a challenge to ensure that the mutated code is not only syntactically correct but also aligns well with the realistic, common coding patterns and practices observed among developers. (3) Forming diverse and killable mutations. Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper explores whether Large-Language Models (LLMs), trained on big code and can produce human-like code (Chen et al., 2021), could deal with these challenges. Following existing studies (Guo et al., 2024; Ma et al., 2024; Schäfer et al., 2023), we designed a default prompt that contains task instructions, the code element to be mutated, the corresponding method code, and a few-shot examples sampled from real-world bugs. We carried out a large-scale evaluation involving six LLMs (three closed-source commercial and three open-source ones), 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)), and compared with existing mutation testing techniques and tools.
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of cost, usability (i.e., compilability rate, useless mutation rate, and equivalent mutation rate, AST node diversity), and behavior similarity with real bugs (i.e., real bug detection rate, coupling rate, and Ochiai Coefficient). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs. Overall, our experiments generate about 7.8 million mutations and execute the origin test suites 5.7 million times.
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o lead to a fault detection rate of 93.4%, while the mutants generated by LEAM, PIT, and Major are 71.7%, 51.3%, and 74.4%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, Major excels with a 98.33% compilability rate, 0% useless mutations, and 0.59% equivalent mutants, while GPT4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%, respectively.
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 57 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 31. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations.
We also analyzed the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall in 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
We wrapped up our LLM-based method into a mutation tool called Wukong. Based on the experimental data, we have also built a dataset that contains high-quality LLM-generated mutants that can be used not only in mutation testing but also in other bug-seeding applications, such as fault localization and fault prediction. The implementation of Wukong and the experimental data are available for open science (our, 2024).
In summary, our paper makes the following main contributions:
  • We investigate the applicability of LLMs in mutation testing. We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that GPT-4 models excel in generating mutations that closely mimic real bugs.
  • We compare different prompts, and find that few-shot learning with suitable code context archives best performance.
  • We analyze the error types of non-compilable mutations and figure out that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
  • We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
2. Background and Related Work
Mutation Testing and Mutation Generation.  Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one popular mutation operator is to change one arithmetic binary operator with another one (e.g., $a+b\mapsto a-b$). The majority of existing mutation testing approaches are rule-based, such as PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanch (Schuler and Zeller, 2009), Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), AccMut (Wang et al., 2017), WinMut (Wang et al., 2021), etc. Then researchers proposed using deep learning techniques to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed $\mu$Bert (Degiovanni and Papadakis, 2022). $\mu$Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches.
LLMs for Code.  Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo, GPT-4o, and GPT-4o-Mini are closed-source LLMs released by OpenAI (Achiam et al., 2023), while DeepSeek-Coder, StarChat, and CodeLlama are open-source. An open-source LLM is usually fine-tuned against a base model for specified tasks. DeepSeek-Coder (Zhu et al., 2024) is based on DeepSeek. CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets. StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
3. Study Design
3.1. Overview and Research Questions
Our paper intends to thoroughly investigate the capability of existing LLMs in mutation generation. We design research questions covering the evaluation of the default setting of our study, the impact of prompt engineering strategies and models, and root cause analysis for underperforming aspects. Specifically, we aim to answer the following five research questions.
  • RQ1: How do the LLMs perform in generating mutations in terms of cost and usability?
  • RQ2: How do the LLMs perform in generating mutations in terms of behavior similarity with real bugs?
  • RQ3: How do different LLMs affect the performance in generating mutations?
  • RQ4: How do different prompts affect the performance in generating mutations?
  • RQ5: What are the root causes and compilation error types for the generation of non-compilable mutations?
3.2. Datasets
Table 1. Real Bugs Used in Our Experiment
Dataset Project # of Bugs Time Span
Defects4J (D4J) Math 106 2006/06/05 - 2013/08/31
Lang 65 2006/07/16 - 2013/07/07
Chart 26 2007/07/06 - 2010/02/09
Time 27 2010/10/27 - 2013/12/02
Closure 133 2009/11/12 - 2013/10/23
Mockito 38 2009/06/20 - 2015/05/20
Cli 39 2007/05/15 - 2018/02/26
Codec 18 2008/04/27 - 2017/03/26
Csv 16 2012/03/27 - 2018/05/18
Gson 18 2010/11/02 - 2017/09/21
JacksonCore 26 2013/08/28 - 2019/04/05
Jsoup 93 2011/07/02 - 2019/07/04
ConDefects (CD) 246 2024/03/01 - 2024/06/30
Total 851 2006/07/16 - 2024/06/30
We intend to evaluate our approach with real bugs, and thus we need to use bug datasets with the following properties:
  • The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
  • The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
  • Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
  • Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 1. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023a; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 2 and Table 1, we observe the time spans of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder [^1] https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
3.3. Mutation Generation via LLMs
Given the location of a Java program, we extract essential information (e.g., the context method or the corresponding unit tests) to formulate prompts instructing mutation generation, then feed prompts to the selected LLM. Once the LLM returns its responses, we filter out mutations that are non-compilable, redundant, or identical to the original code, and return the generated mutations. As shown in Section 3.4, our study supports a comprehensive set of metrics for evaluating mutations.
3.3.1. Studied Models
We aim to comprehensively compare the existing LLMs in their ability to generate mutations from various perspectives. This includes evaluating the capabilities of models that are fine-tuned on different base models and contrasting the performance of commercial closed-source models with open-source alternatives. In our study, we select a representative set of LLMs with code capabilities. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from GPT family, including GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-Mini (Achiam et al., 2023). Moreover, we employ the popular open-sourced LLMs that are used by recent studies (Ni et al., [n. d.]; Xia et al., 2024; Ma et al., 2024; Yang et al., 2024), including DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023). The details of these models are shown as Table 2. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
Table 2. Studied LLMs
Model Type Studied Model Base Model Training Data Time Release Time Size
Closed GPT-3.5-Turbo GPT 2021/09 2023/03
GPT-4o GPT 2023/10 2024/05
GPT-4o-Mini GPT 2023/10 2024/07
Open StarChat-$\beta$-16b StarCoder 2023/06 16B
CodeLlama-Instruct-13b Llama 2023/08 13B
DeepSeek-Coder-V2-236b DeepSeek 2023/09 2024/07 236B
3.3.2. Prompts
Prompts are crucial for LLMs because they guide the models’ responses and determine the direction of the generated output. A well-crafted prompt can significantly influence the utility, relevance, and accuracy of the code produced by an LLM. This subsection introduces how we design the prompts.
Figure 1 shows the default prompt template in our study. To design effective prompts for mutation generation, we follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our default prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
Table 3. Default Few-Shot Examples from QuixBugs
Correct Version Buggy Version
n = (n & (n - 1)); n = (n ^ (n - 1));
while (!queue.isEmpty()) while (true)
return depth==0; return true;
ArrayList r = new ArrayList();
r.add(first).addll(subset);
to_add(r);
to_add.addAll(subset);
c = bin_op.apply(b,a); c = bin_op.apply(a,b);
while(Math.abs(x-approx*approx)>epsilon) while(Math.abs(x-approx)>epsilon)
Figure 1. The Default Prompt Template
3.4. Evaluation Metrics
To assess the utility of the LLM-generated mutations, we should define a set of metrics that comprehensively evaluate them. Therefore, we designed several categories of metrics capturing different aspects as follows: Cost Metric, Usability Metrics, and Behavior Metrics. Besides these metrics, we also additionally report the Mutation Count and Mutation Score. Note that both of them cannot directly reflect the quality of generated mutations. For example, a higher mutation score does not indicate a higher quality, which may also be caused by generating a large proportion of mutants that are easily killed.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set $A$, where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set $C$. Note that the set $C$ contains two types of mutations that should be removed, the useless mutations (denoted as $U$) and equivalent mutations (denoted as $E$). The useless mutations refer to those that are identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set $(C-E-U)$, excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set $E$ is undecidable, so we follow the common practice and use the set $(C-U)$ as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022).
3.4.1. Cost Metric
The Cost Metric indicates time costs in mutation generation.
Average Generation Time:  the average duration needed to produce each mutation, measuring the efficiency of an approach.
3.4.2. Usability Metrics
The Usability Metrics serve as key indicators of the practical utility of generated mutations.
Compilability Rate:  the proportion of mutations that successfully compile, computed as: ${|C|}/{|A|}$. Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
Useless Mutation Rate:  the proportion of useless mutations, computed as: $|U|/|A|$. LLMs sometimes generate duplicate mutations or mutations that are exactly the same as the original code. While traditional approaches rarely generate useless mutations, LLMs suffer from issues of misunderstanding and repetition, leading to generating these useless mutations.
Equivalent Mutation Rate:  the proportion of equivalent mutation, computed as: $|E|/|A|$. Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. Due to the undecidable nature of identifying whether a mutation is equivalent, we sample significant portions of all mutations and judge them manually. More specifically, given the population to be the set $A$, we set the confidence level to be 95% and the margin of error to be 5%, to get a subset of mutations to be manually inspected.
AST Node Diversity:  We prefer the set of mutations with diverse forms, therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior.
3.4.3. Behavior Similarity with Real Bugs
Behavior similarity refers to the degree of similarity in functionality and execution between a mutation and its corresponding real bug. Because capturing all program behaviors is undecidable, following existing studies, we thus employ the following metrics as a proxy to measure behavior similarity (Ojdanic et al., 2023a; Papadakis et al., 2018; Khanfir et al., 2023).
Real Bug Detectability:  Given the location of a real bug (which may occupy multiple lines), let a mutation generation approach generate a set of mutations, if the tests that kill these mutations can also kill this real bug, we say that the mutation generation approach can “detect” this real bug. Given a set of real bugs and a mutation generation approach, Real Bug Detectability quantifies the number of bugs detected by this mutation approach.
Coupling Rate:  Mutation testing is based on the fault coupling effect, i.e., mutations though are simple syntactic changes, coupled to more complex bugs (Offutt, 1992). To describe the extent to which a mutation couples to its corresponding real bug, we adopt the Coupling Rate following the existing study (Just et al., 2014b). Given a real bug and a corresponding mutant generated at the same location, the mutant is deemed a coupled mutant if the set of the test cases killing the mutant overlaps with the set of the bug-triggering test cases. The Coupling Rate refers to the proportion of coupled mutants within $(C-U)$.
Ochiai Coefficient:  approximates the semantic similarity by passing and failing test cases (Offutt and Untch, 2001), which is also widely used in the field of fault localization. Following the existing study (Ojdanic et al., 2023a), we employ the Ochiai Coefficient to calculate how a mutation resembles its corresponding real bug. More concretely, let $M$ and $B$ be a mutation and its corresponding real bug, and let $fT_{M}$ and $fT_{B}$ be the set of their failing test cases respectively. $M$’ coefficient is computed by the following formula:
(1) $Ochiai(M,B)=\frac{|fT_{M}\cap fT_{B}|}{\sqrt{|fT_{M}|\times|fT_{B}|}}$
Therefore, if $M$’s Ochiai Coefficient is $1$, it is semantically equal to its corresponding bug $B$.
3.5. Experiment Settings
3.5.1. Settings for Mutation Generation
One major target of our study is to investigate the similarity between LLM-generated mutations and real bugs. Therefore, we generate mutations within the context of real bugs to enable such a comparison. To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the full set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 5.
3.5.2. Settings for RQ1-RQ3
In the first part of our study, we intend to qualitatively and comparatively evaluate the capability of LLMs in generating mutations. We adopt the default settings, which contain all six LLMs, and use the default prompt template, generating mutations on the universal set of the 851 bugs in Section 3.2.
To understand the performance of the LLMs, we adopt LEAM (Tian et al., 2022), $\mu$Bert (Degiovanni and Papadakis, 2022), PIT (Coles et al., 2016), and Major (Just et al., 2011) as baselines. Among these approaches, LEAM is the state-of-the-art mutation generation approach which employs a model trained from 13 million real Java bugs. While $\mu$Bert is based on Bert (Devlin et al., 2018), a masked pre-trained language model that can not conversational respond to human-like text and code.PIT (Coles et al., 2016) and Major (Just et al., 2011) are the popular traditional mutation testing tools with human-predefined operators. Note that for PIT and Major, we employ all mutation operators.
3.5.3. Settings for RQ4
In the second part of our study, we intend to explore how different prompts and different models impact the performance of mutation generation. Due to budget limitations, we sample a subset of 306 bugs from our dataset, which consists of 10 sampled bugs from each Defects4J project and all ConDefects bugs.
In RQ4, for the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get 4 types of prompts, shown as follows:
  1. (1)
    Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
  2. (2)
    Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the examples.
  3. (3)
    Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
  4. (4)
    Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
3.5.4. Settings for RQ5
Generating non-compilable mutations incurs additional meaningless economic and time costs, so in the third part of our study, we study the root causes behind LLM generating non-compilable mutations. Specifically, we sample 384 non-compilable mutations of each mutation generation approach and analyze their reason for being rejected by the Java compiler and the features of their surrounding code context.
4. Evaluation Results
We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
4.1. RQ1: Performance on Cost and Usability
4.1.1. Cost
For each technique, we record the generation duration, as shown in the Avg. Gen. Time row of Table 4. In addition, we report the mutation count and mutation score, although they do not fully reflect the overall quality of the mutation set.
From the table, we find that GPT-3.5, GPT-4o, GPT-4o-Mini, DeepSeek, StarChat, and CodeLlama generate 538338, 539715, 588803, 594711, 463248, and 457195 mutations, respectively222 Although the LLMs are instructed to generate the same number of mutations, they do not consistently follow the instructions, thereby yielding different numbers of mutations., with a mutation score ranging from 0.516 to 0.726. Overall, traditional approaches (e.g., LEAM and PIT) yield lower mutation scores among all the approaches.
In terms of time cost, the six LLMs cost 1.79s, 1.76s, 1.65s, 4.25s, 7.53s, and 9.06s to generate mutations, respectively. In comparison, LEAM, $\mu$Bert, PIT, and Major take 3.06s, 2.23s, 0.02s, and 0.08s, respectively. The rule-based approaches, PIT and Major, are the fastest, while LLMs show a competitive speed with small model-based approaches, LEAM and $\mu$Bert.
Finding 1 (RQ1): Rule-based approaches cost less time than LLMs in generating mutations. For example, GPT LLMs take fewer than 1.8s per mutation. In contrast, PIT and Major only require 0.02s and 0.08s, respectively, while small model-based approaches (i.e., LEAM and $\mu$Bert) take 2-3s.
Table 4. Overall Performance of All the Mutation Generation Techniques
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
Mut. Score 0.697 0.726 0.708 0.694 0.516 0.700 0.545 0.675 0.581 0.486
Mut. Count 538338 539715 588803 594711 463248 457195 890417 200896 2890433 346282
Avg. Gen. Time 1.79 1.76 1.65 4.25 7.53 9.06 3.06 2.33 0.02 0.08
Comp. Rate 60.2% 75.6% 73.6% 75.5% 11.1% 70.2% 35.0% 22.5% —— 98.6%
Useless Mut. Rate 10.9% 7.8% 6.7% 8.3% 8.9% 39.3% 1.0% 1.7% 0.0% 0.0%
Eq. Mut. Rate 2.2% 1.2% 1.7% 1.2% 1.8% 1.2% 1.3% 2.5% 0.0% 0.6%
Real Bug Detec. 91.7% 93.4% 93.4% 92.8% 47.1% 83.1% 71.7% 71.3% 51.3% 74.4%
Coupling Rate 41.4% 43.6% 44.1% 41.5% 29.4% 39.8% 28.7% 41.6% 14.0% 36.0%
Ochiai Coeff. 65.0% 68.5% 66.9% 61.8% 19.3% 40.8% 37.4% 31.6% 31.5% 44.4%
4.1.2. Mutation Usability
The usability of mutations (i.e., whether the mutations can be used to calculate mutation score) is affected by the capabilities of LLMs in correctly understanding the task and generating compilable, non-duplicate, and non-equivalent mutations. The results are shown as the Comp. Rate, Useless Mut. Rate, and Eq. Mut. Rate rows in Table 4. Since the mutation generation process of the PIT model is not visible and it only outputs mutations that pass the tests, we excluded it from the compilability rate calculation.
Compilability Rate: Four out of the six LLMs (i.e., GPT-4o, GPT-4o-Mini, DeepSeek, CodeLlama) achieve over 70%, with one exceeding 60% (i.e., GPT-3.5). In contrast, the small model-based approaches, LEAM and $\mu$Bert, have rates at 35.0% and 22.5%, respectively. Compared to small models, LLMs like GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 40% higher than that of LEAM and $\mu$Bert. The rule-based approach Major significantly outperforms, with nearly 99% of its mutations successfully compiling. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true).
Useless Mutation Rate: We also observe that LLMs generate a significant portion of duplicate and useless mutations. For useless mutations, the percentage generated by LLMs ranges from a minimum of 6.7% (GPT-4o-Mini) to a maximum of 39.3% (CodeLlama), and five out of the six LLMs have a useless mutation rate below 11%. In contrast, all the traditional approaches rarely generate duplicate mutations, which is likely due to them being governed by strict grammar rules that limit the production of redundant mutations. For example, PIT and Major iteratively apply different mutation operators producing no redundant mutants.
Table 5. Sample Results for Equivalent Mutations
Dataset Type GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
D4J #All 537109 538509 587583 593481 462042 455634 888452 199347 2888641 345292
#Sampled 384 384 384 384 384 384 384 384 384 384
#Eq. Mut 9 3 5 4 8 5 6 8 0 2
CD #All 1229 1206 1220 1230 1206 1561 1965 1549 1792 990
#Sampled 293 292 293 293 292 309 322 308 317 277
#Eq. Mut 6 5 7 4 4 3 3 9 0 2
Equivalent Mutation Rate: Equivalent mutations change the source code syntactically without affecting the program’s functionality, thereby affecting the accuracy of mutation score calculations. Due to the undecidable nature of identifying equivalent mutations (Budd and Angluin, 1982), it is not feasible to design an algorithm that can completely detect them. Therefore, we sample a subset of mutations for each technique and manually assess whether they are equivalent. Following the common practice of sampling (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015), to maintain a 95% confidence level and a 5% margin of error, we sampled a fixed number of mutations from each approach’s output, to calculate the Equivalent Mutation Rate. Following the protocol of existing studies (Islam et al., 2019; Guo et al., 2024), two authors, each with over five years of Java programming experience, first received training on the equivalent mutants, and then independently labeled the mutations. The final Cohen’s Kappa coefficient score (Viera et al., 2005) is more than 90%, indicating a high level of agreement. All labeling results are available in our package for reproducibility (our, 2024), and Table 5 presents the sampling and labeling results.
As shown in the Eq. Mut. Rate row of Table 4, we again observe that the rule-based approaches, PIT and Major, perform the best in generating non-equivalent mutants, with an equivalent mutation rate of less than 1.0%. All six LLMs exhibit an equivalent mutation rate ranging from 1.2% to 2.2%, similar to LEAM and $\mu$Bert. Overall, the equivalent mutation rates for all approaches studied are relatively low and can be ignored during mutation testing. We further analyzed the mutations generated by PIT and Major, and found that they apply specific mutation operators to a statement (e.g., for a < b, only generating !(a < b), true, and false), significantly reducing the equivalent mutation rate.
Finding 2 (RQ1): Rule-based approaches, PIT and Major, outperform others in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, GPT-4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%. In contrast, Major excels with a compilability rate of 98.3%, a useless mutation rate of 0%, and an equivalent mutation rate of 0.6%.
Figure 2. Number of newly introduced AST nodes of the studied approaches and the few-shot examples.
4.1.3. AST Node Dirversity
To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 shows the number of introduced types, which reveals that LLM-based mutation approaches introduce more types of new AST nodes than other approaches. The best of LLMs, GPT-4o-Mini and DeepSeek, exhibit the greatest diversity, introducing 57 new AST node types. Among the traditional approaches, LEAM, $\mu$Bert, and Major introduce 31, 30, and 2 new types of AST nodes, respectively. Additionally, all six LLMs significantly are higher than traditional approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
Table 6. Ratio of Deletion Mutations and the Top-3 Newly Introduced AST Node Types for Each Approach.
Top K GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major
Del. 5.7% 2.2% 0.5% 1.1% 0.2% 0.1% 48.4% 0.6% 20.2%
1st LT(36.1%) LT(43.5%) LT(41.7%) LT(40.1%) LT(20.1%) LT(42.5%) LT(21.4%) LT(25.4%) ST(58.6%)
2nd BO(17.4%) BO(16.5%) BO(22.0%) BO(18.7%) MR(13.7%) BO(23.8%) MR(14.7%) SE(21.7%) LT(41.4%)
3rd MI(10.9%) MI(7.9%) MI(7.1%) MI(10.5%) BO(11.8%) MI(8.0%) MI(12.0%) MR(21.9%) ——
LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 6, which highlights the top three most frequent AST node types. For the proportion of deletions, LEAM shows the highest deletion mutation ratio, with 48.4%, followed by Major at 20.2%. In contrast, all six LLMs generate deletion mutations at a rate below 6%, with three models producing fewer than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we found that deletions are one of the key factors contributing to the lower compilation success rate of traditional approaches.
Regarding AST node types, the most frequent new AST node type of all the learning-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the rule-based approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
Finding 3 (RQ1): LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches, while less inclined to generation deletion mutations.
4.2. RQ2: Behavior Similarity
The bottom three rows of Table 4 present a comparative evaluation of Behavior Metrics for mutation generation approaches.
4.2.1. Real Bug Detectability
Among the six LLMs, five models (excluding StarChat) outperform all the traditional approaches. Of the traditional approaches, the rule-based approach Major performs the best, detecting 74.4% of the bugs. GPT-4o and GPT-4o-Mini achieve the highest detection rates, identifying 93.4% of the 851 bugs, excelling Major by 19% on all bugs.
In particular, on the ConDefects dataset, which mitigates data leakage risks, the real bug detection rates for the various approaches are as follows: GPT-3.5 (78.9%), GPT-4o (84.6%), GPT-4o Mini (82.5%), DeepSeek (83.7%), StarChat (32.9%), CodeLlama (63.4%), LEAM (48.9%), $\mu$Bert (47.2%), PIT (49.2%), and Major (56.1%). The best of LLMs (i.e., GPT-4o) outperforms nearly 29% over the best of traditional approaches (i.e., Major), highlighting the significant advantage that LLMs hold over traditional techniques in detecting real bugs.
Finding 4 (RQ2): In Real Bug Detectability, five out of six LLMs significantly outperform traditional approaches. In particular, on ConDefects, GPT-4o outperforms nearly 29% over the best of the traditional approach, Major, highlighting the advantages of LLMs in detecting real bugs.
4.2.2. Coupling Rate
The Coupling Rate measures the degree of coupling between the generated mutations and their corresponding real bugs.
Among traditional approaches, $\mu$Bert is the best one with a rate of 41.6%. However, two LLMs outperform all traditional approaches in this metric, achieved by the GPT-4 LLMs. The best LLM, GPT-4o-Mini, exhibits a rate of 44.1%, exceeding $\mu$Bert by 2.5%.
Finding 5 (RQ2): GPT-4 achieves the highest Coupling Rate at 44.1%, outperforming all traditional approaches, including the best traditional approach, $\mu$Bert, by 2.5%.
4.2.3. Ochiai Coefficient
The Ochiai Coefficient measures the semantic similarity between mutations and real bugs (Ojdanic et al., 2023a).
Similar to real bug detectability, most LLMs significantly outperform all existing approaches. Among the LLMs, GPT-4o achieves the highest Ochiai coefficient at 68.5%, 24.1% higher than the best rule-based approach, Major (44.4%), and 31.1% higher than the best small model-based approach, LEAM (37.4%). Moreover, five out of the six LLMs (excluding StarChat), outperforms all traditional approaches.
Finding 6 (RQ2): In the Ochiai Coefficient, five LLMs surpass all traditional generation approaches. In particular, the best LLM, GPT-4o, exceeds the best of traditional approaches, Major, by 24.1%.
4.3. RQ3: Impacts of Different LLMs
As shown in Table 4, we compare the six LLMs we studied in mutation generation. The results reveal significant performance differences among the models.
For Cost, GPT-4o-Mini is the fastest, requiring only 39% of the time taken by the fastest open-source model, DeepSeek.
For Usability, GPT-4o and GPT-4o-Mini lead in all metrics with similar performance, while DeepSeek closely follows and performs similarly to GPT-3.5. Additionally, as shown in Figure 2, GPT-4o-Mini and DeepSeek exhibit the most diverse AST node types, indicating their mutations are more expressive.
For Behavior, GPT-4o and GPT-4o-Mini excel in all metrics, and DeepSeek has similar performance to GPT-3.5, making it the best open-source LLM. StarChat and CodeLlama underperform across most metrics, particularly in compilation rate and bug detection. Based on the model information in Table 2, we find that newer models tend to generate higher-quality mutations, which aligns with the trends observed in other code-related tasks (Yang et al., 2024).
Finding 7 (RQ3): Two GPT-4 models perform best, while DeepSeek has a similar performance to GPT-3.5, standing out among open-source models. The newer models exhibit better performance in mutation generation.
4.4. RQ4: Impacts of Different Prompts
Table 7. Comparison Results of Different Prompts
Metric Category Metric GPT-3.5 DC-236b
P1 P2 P3 P4 P1 P2 P3 P4
—— Mut. Score 0.708 0.706 0.736 0.782 0.712 0.723 0.719 0.737
Mut. Count 48670 49126 42775 40844 53000 53177 52687 53130
Cost Avg. Gen. Time 1.79 1.78 1.80 1.87 4.25 4.25 4.20 4.40
Usability Compilability Rate 65.7% 63.8% 61.5% 60.6% 74.2% 75.1% 70.2% 72.6%
Useless Mut. Ratio 13.6% 11.3% 21.8% 14.8% 8.9% 3.8% 11.7% 9.8%
Eq. Mut. Rate 2.2% 1.9% 2.4% 2.5% 1.2% 1.3% 1.0% 1.5%
Behavior Real Bug Detectability 81.4% 80.7% 75.5% 68.0% 85.0% 83.7% 82.0% 82.7%
Coupling Rate 15.4% 15.1% 11.6% 11.7% 14.8% 14.6% 13.9% 14.4%
Ochiai Coefficient 66.3% 65.4% 63.1% 64.7% 66.7% 64.1% 64.4% 65.0%
Table 7 presents the comparative results of GPT-3.5 and DeepSeek using different prompts listed in Section 3.5.3. Prompts P1 to P3 simplify progressively, each containing less information than its predecessor, while P4 is the most complex, enhancing P1 with test suite code.
Overall, P1, the default prompt, excels in all Behavior Metrics on both models, indicating it can generate mutations more likely to resemble real bugs and effectively detect them. In addition, for all Usability Metrics, P1 achieves the second-best performance. P2, created by removing few-shot examples from P1, performs better in usability. P3, provided only with the code element to be mutated, achieves the lowest equivalent mutation rate. Conversely, P4, which extends P1 with additional test suite code, performs the worst in most cases.
Finding 8 (RQ4): The prompt with the whole method and few-shot examples as context (i.e., P1) achieves the best performance across all Behavior Metrics, whereas adding the code of test suites (i.e., P4) decreases performance.
4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 4.1, LLMs generate a significant number of non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT-3.5, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified 9 distinct error types, as shown in Table 8.
Table 8. Error Types of Non-Compilable Mutations
ID Error Type % ID Error Type %
1 Usage of Unknown Methods 27.3% 6 Type Mismatch 7.6%
2 Code Structural Destruction 22.9% 7 Incorrect Initialization 3.1%
3 Incorrect Method Parameters 12.8% 8 Incorrect Location 3.1%
4 Usage of Unknown Variables 11.2% 9 Incorrect Exceptions 2.3%
5 Usage of Unknown Types 9.6%
Shown as Table 8, the most common error type, Usage of Unknown Methods, accounts for 27.3% of the total errors, revealing the hallucination issue of generative models (Huang et al., 2023). Code Structural Destruction is the second most common error, accounting for 22.9%, indicating that ensuring generated code is syntactically correct remains a challenge for current LLMs. This result suggests that there is still significant room for improvement in current LLMs.
Finding 9 (RQ5): The mutations generated by GPT models have nine main types of compilation errors, with Usage of Unknown Methods and Code Structural Destruction being the two most prevalent types.
(a) GPT-3.5
(b) GPT-4o
(c) GPT-4o-M
(d) DS-236b
(e) SC-16b
(f) CL-13b
(g) LEAM
(h) $\mu$Bert
(i) Major
Figure 3. Distribution of the AST Node Types of Origin Code for Non-Compilable Mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types. For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% ($\mu$Bert), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation. Thus for LLMs, deletion is not the major reason for non-compilation.
Finding 10 (RQ5): Member references and method invocations within the code context are the most likely triggers for LLMs to generate non-compilable mutations.
5. Discussion
5.1. Sensitivity to Chosen Experiment Settings
5.1.1. Experiment Setup of the Context Length
Table 9. Performance of Different Context Lengths
Metric Category Metric 1-Line Context 2-Line Context 3-Line Context
Usability Compilability Rate 0.610 0.617 0.657
Useless Mutation Ratio 0.119 0.128 0.137
Equivalent Mutation Rate 1.6% 1.1% 1.3%
Behavior Real Bug Detectability 0.933 0.917 0.917
Coupling Rate 0.141 0.121 0.136
Ochiai Coefficient 0.517 0.417 0.467
Table 10. Performance Similarity of Different Context Lengths
Settings Spea. Coef. $p$-value Pear. Coef. $p$-value
1-Line v.s. 2-Line 0.9762 3.31E-05 0.9980 2.12E-08
2-Line v.s. 3-Line 0.9762 3.31E-05 0.9997 1.04E-10
1-Line v.s. 3-Line 0.9524 2.60E-04 0.9978 2.69E-08
QB-6a (Default) v.s. QB-3 0.9701 6.55E-05 0.9996 1.70E-10
QB-6a (Default) v.s. QB-3b 0.9762 3.31E-05 0.9984 1.00E-08
QB-6a (Default) v.s QB-9 0.9762 3.31E-05 0.9985 7.98E-09
QB-6a (Default) v.s D4J-6 0.9524 2.60E-04 0.9980 2.13E-08
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 9. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the upper part of Table 10. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
5.1.2. Impact of Different Few-Shot Examples
Table 11. Impact on Different Few-Shot Examples
Metric QB-6a (Default) QB-3 QB-6b QB-9 D4J-6
Comp. Rate 65.7% 62.7% 63.9% 61.8% 62.1%
Useless Mut. Ratio 13.7% 12.8% 12.8% 13.8% 12.5%
Equivalent Mutation Rate 1.3% 1.6% 1.3% 1.1% 1.6%
Real Bug Det. 0.917 0.933 0.917 0.917 0.900
Coupling Rate 0.136 0.128 0.132 0.121 0.139
Ochiai Coefficient 0.467 0.450 0.434 0.434 0.367
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 11.
We further compute their similarity, shown as the lower part of Table 10. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude the performance of different prompts is very similar.
5.1.3. Performance of Using the Same Number of Mutations
In Section 4, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 16,785. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 12. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
Table 12. Performance Under the Same Number of Mutations
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major PIT
Comp. Rate 62.9% 77.4% 70.1% 73.0% 39.3% 74.2% 38.9% 18.5% 96.5% ——
Useless Mut. Rate 11.0% 7.3% 7.0% 8.9% 27.1% 39.0% 1.4% 1.4% 0.0% ——
Real Bug Det. 89.2% 90.8% 91.3% 89.7% 58.0% 73.5% 67.5% 55.0% 83.3% 43.7%
Coupling Rate 13.3% 14.6% 12.6% 13.4% 11.9% 10.5% 13.2% 13.2% 12.8% 2.4%
Ochiai Coefficient 35.8% 45.0% 40.8% 35.3% 15.3% 17.8% 17.0% 15.0% 34.1% 19.2%
5.2. Economic Costs
Closed-source LLMs (e.g., GPT LLMs) are typically billed based on the number of tokens used. As these models continue to evolve, their pricing rates also change, and we here discuss the costs of the proprietary models used in our experiments. For the three models involved, GPT-3.5 incurred a total cost of \$108, GPT-4o cost \$1531, and GPT-4o-Mini cost \$47. Therefore, for every 1,000 mutations, GPT-3.5, GPT-4o, and GPT-4o-Mini cost about \$0.16, \$2.84, and \$0.08, respectively. GPT-4o is significantly more expansive than the others. Given the effectiveness of mutations generated by GPTs, the cost is feasible for large-scale applications.
5.3. Implications
Based on our findings, we discuss the pros and cons of using LLMs for generating mutations and suggest potential improvements.
First, our study shows that existing LLMs have the potential to generate high-quality mutations for mutation testing with acceptable costs. The greatest advantage of LLMs is their ability to generate mutations that more closely mimic real bugs in behavior. For example, Listing 1 shows two compilable mutations generated by GPT-3.5 in Chart-1, which are theoretically beyond the capabilities of existing approaches. In Mutation-1, the LLM inferences there must exist leftBlock because of the occurrence of rightBlock, and replaces the variable correspondingly. This mutation is not killed by the bug-triggering test. Without natural language understanding capabilities (i.e., understanding the correspondence between “right” and “left”), it is difficult to create such mutations. In Mutation-2, the LLM modifies the else branch to else if, employing an integer variable seriesCount from context to synthesize the condition, showing the capabilities of code structural changing that existing methods cannot achieve. Also, this mutation is alive and can guide test enhancement.
// Mutation-1
-   Size2D size = this.rightBlock.arrange(g2, c4);
+   Size2D size = this.leftBlock.arrange(g2, c4);
// Mutation-2
-   else {
-     g2.setStroke(getItemOutlineStroke(row, column, selected));
+   else if (seriesCount == 1) {
+     g2.setStroke(getItemOutlineStroke(row, column, selected));
Listing 1: Two Example Mutations of GPT-3.5 on Chart-1
Second, our study explores how different prompts, models, contexts, and few-shot examples impact mutation generation performance. Our study highlights the importance of using proper prompts and models with stronger coding capabilities. In particular, designing prompts requires providing essential context information (such as the whole method), but including too much information (such as the target unit test) may reduce effectiveness.
Third, although LLMs achieve promising results in mutation testing, our study indicates that there is still significant room for improvement. Specifically, current LLMs generate a substantial number of mutations that are not compilable. Therefore, we analyze the types of mutations that lead to non-compilable code and identify some root causes of the LLMs. We advocate for more research efforts in program analysis and repair approaches to address error-prone code elements for mutation such as method invocations.
5.4. Threats to Validity
The selected LLMs, programming language, datasets, and baseline approaches could be a threat to the validity of our results. To mitigate this threat, we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek), the most popular language (i.e., Java), and the most popular dataset Defects4J. We also employ state-of-the-art mutation testing approaches as baselines, including learning-based and rule-based.
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (469), GPT-4o (290), GPT-4o-Mini (528), DeepSeek (318), StarChat (122), CodeLlama (236), LEAM (590), $\mu$Bert (66), and Major (217), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), $\mu$Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of the exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar.
The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
6. Conclusion
In this paper, we systematically investigate the performance of LLMs in generating mutations. We evaluate their utility from several aspects and find LLMs have the advantage of generating diverse mutations that mimic the behaviour of real bugs. We also analyze and discuss some limitations of LLMs and point out further directions. We advocate research efforts to promote LLM for mutation testing in the future.
7. Data Availability
All the experimental data, code, and full results (including generated and manually labeled mutations) are publicly available (our, 2024).