### v3

# A Comprehensive Study on Large Language Models for Mutation Testing

Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we report the results of a comprehensive empirical study over six different LLMs, including both state-of-the-art open- and closed-source models, on 851 real bugs drawn from two different Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with $90.1\%$ higher fault detection. That is, 79.1% (for LLMs) vs. 41.6% (for rule-based); an increase of $37.5$ percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by $36.1$, $13.1$, and $4.2$ percentage points respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
Empirical Study, Mutation Testing, Mutation Analysis, Mutation Generation, Large Language Models, Coupling
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debugging
1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In those most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Foster et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021) reported on results from deployment of mutation testing at their companies. In particular, results recently reported by Meta (Foster et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, there is to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique is challenging because of the following three important issues:
  1. (1)
    Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
  2. (2)
    Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed. Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
  3. (3)
    Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper reports on how well Large-Language Models (LLMs) handle these challenges, compared to more traditional rule-based approaches. We conducted a comprehensive comparison between six LLMs (three closed-source commercial and three open-source ones) and four existing mutation testing techniques with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our experiments required the generation of 611,100 mutations.
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to a higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o-Mini lead to a fault detection rate of 90.8%, while the mutants generated by LEAM, PIT, and Major are 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also present that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have the AST edit distance of one, while for GPT-4o, only 50% of its mutations have the same distance.
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
In summary, our paper makes the following main contributions:
  • We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
  • We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
  • We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
The remainder of this paper is organized as follows. Section 2 presents the design of our study, including the research questions, studied approaches, datasets, and metrics. Section 3 presents the empirical evaluation of these generation approaches and answers to each research question. Section 4 discusses the influences of different experimental setups and the implications summarized from the findings. Section 5 presents the threats to the validity of our study. Section 6 discusses the related work. Section 7 concludes this work.
2. Study Design
In this section, we introduce how our study is designed and conducted.
2.1. Overview and Research Questions
This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
  • RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutations in representing real bugs. Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
    • RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
    • RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
    • RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
  • RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of mutations generated by LLMs compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
    • RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
    • RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
    • RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
  • RQ3 (Efficiency) What are the time costs of different mutation generation approaches? This research question investigates the time costs of different mutation generation approaches. Time costs are associated with the efficiency and practicality of mutation generation approaches.
  • RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches. Understanding these errors can guide mutation generation approaches to enhance their practical usability.
2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking [^1] https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches is for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. Table 1 summarizes the mutation generation approaches covered by this study.
Table 1. Studied Mutation Generation Approaches
Type Name Description
Traditional PIT Widely studied mutation testing framework with 29 mutation operators.
Major Widely studied mutation testing framework with 11 mutation operators.
LEAM Custom-trained models learning from a large corpus of real bugs and patches.
$\mu$Bert Mutation generation with the pre-trained model CodeBERT.
LLM-based GPT-3.5 GPT-3.5-Turbo, whose based model is GPT, trained on data before 2021/09, released in 2021/03.
GPT-4o The based model is GPT, trained on data before 2023/10, released in 2024/05.
GPT-4o-M GPT-4o-Mini, whose based model is GPT, trained on data before 2023/10, released in 2024/07.
SC-16b StarChat-$\beta$-16b, whose based model is StarCoder, released in 2023/06.
CL-13b CodeLlama-Instruct-13b, whose based model is Llama, released in 2023/08.
DS-236b DeepSeek-Coder-V2-236b, whose based model is DeepSeek, released in 2024/07.
2.2.1. Traditional Approaches
For Java, there exist several mutation testing frameworks, such as JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), PIT (Coles et al., 2016), and Major (Just et al., 2011). As some tools are no longer maintained, following existing studies (Kaufman et al., 2022; Ojdanic et al., 2023; Kushigian et al., 2024), we employ PIT (Coles et al., 2016), Major (Just et al., 2011), LEAM (Tian et al., 2022), and $\mu$Bert (Degiovanni and Papadakis, 2022). Major and PIT are widely studied rule-based approaches, while LEAM and $\mu$Bert involve deep learning models.
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
$\mu$Bert: $\mu$Bert (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, Bert (Devlin et al., 2018). Equipped with the pre-trained model, $\mu$Bert transforms the task of mutation generation to code completion under the surrounding context.
2.2.2. LLM-based Approaches
Our study involve multiple kinds of LLMs, including both open-source and closed-source models.
LLMs: Our study also comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat-$\beta$-16b), CL-13b (CodeLlama-Instruct-13b), and DS-236b (DeepSeek-Coder-V2-236b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
2.3. Datasets
Table 2. Real-world Bugs Used in Our Study
Dataset Project # of Bugs Time Span
Defects4J (D4J) Math 106 2006/06/05 - 2013/08/31
Lang 65 2006/07/16 - 2013/07/07
Chart 26 2007/07/06 - 2010/02/09
Time 27 2010/10/27 - 2013/12/02
Closure 133 2009/11/12 - 2013/10/23
Mockito 38 2009/06/20 - 2015/05/20
Cli 39 2007/05/15 - 2018/02/26
Codec 18 2008/04/27 - 2017/03/26
Csv 16 2012/03/27 - 2018/05/18
Gson 18 2010/11/02 - 2017/09/21
JacksonCore 26 2013/08/28 - 2019/04/05
Jsoup 93 2011/07/02 - 2019/07/04
ConDefects (CD) 246 2024/03/01 - 2024/06/30
Total 851 2006/07/16 - 2024/06/30
As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
  • The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
  • The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
  • Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
  • Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 2. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 1 and Table 2, we observe the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder [^2] https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives as before mentioned. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set $A$, where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set $C$. Note that the set $C$ contains two types of mutations that should be removed, the duplicate mutations (denoted as $D$) and equivalent mutations (denoted as $E$). Therefore, we have $(C\supseteq D)\wedge(C\supseteq E)\wedge(D\cap E=\emptyset)$. The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set $(C-E-D)$, excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set $E$ is undecidable, so we follow the common practice and use the set $(C-D)$ as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed as bellow.
RQ1.1: Real Bug Detection Rate:  Following the approach of Just et al. (Just et al., 2014b), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
RQ1.2: Average Ochiai Coefficient:  This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al. (Ojdanic et al., 2023).
More concretely, let $m$ and $b$ be a mutation and its corresponding real bug, and let $fT_{m}$ and $fT_{b}$ be the set of their failing test cases respectively. The coefficient of $m$ is computed by the following formula:
(1) $Ochiai(m,b)=\frac{|fT_{m}\cap fT_{b}|}{\sqrt{|fT_{m}|\times|fT_{b}|}}$
Therefore, if the Ochiai coefficient of $m$ is $1$, it is semantically equal to its corresponding bug $b$. Given a bug $b$, its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., $(C_{b}-D_{b})$), shown as the following formula:
(2) $Ochiai_{b}=\frac{\sum_{m\in(C_{b}-D_{b})}Ochiai(m,b)}{|C_{b}-D_{b}|}$
Given a mutation generation approach and a set of bugs $B$, the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
(3) $AOC=\frac{\sum_{b\in B}{Ochiai_{b}}}{|B|}$
RQ1.2: High Ochiai Bug Count:  Following the study (Ojdanic et al., 2023), a bug $b$ and a set of mutations are considered to have high Ochiai if they satisfy $(Ochiai_{b}\geq 0.8)$. This threshold indicates a strong semantic similarity between the generated mutations and the real bug. For each mutation generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate:  Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., $(C-D)$). Note that Real Bug Detection Rate estimates the effectiveness via the proportion of detected bugs, which Coupling Rate estimates it via the proportion of mutations.
RQ1.2: Coupling Categorization:  We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutation (i.e., an element from the set $C$) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
  • Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
  • Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
  • Partial Substitution: Some triggering tests fail and no additional tests fail.
  • Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
  • Not Substitution: Only non-triggering tests fail.
  • Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
Later, we will classify the mutations of each approach and analyze their distributions.
RQ1.3: AST Node Diversity:  We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance:  The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes.
2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate:  the proportion of mutations that successfully compile, is computed as:
(4) $CR=\frac{\enspace|C|\enspace}{|A|}$
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
RQ2.2: Duplicate Mutation Rate:  the proportion of duplicate mutations, computed as:
(5) $DMR=\frac{\enspace|D|\enspace}{|A|}$
LLMs sometimes generate duplicate mutations, i.e., mutations that are exactly the same as the original code. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
RQ2.3: Equivalent Mutation Rate:  the proportion of equivalent mutation, computed as:
(6) $EMR=\frac{\enspace|E|\enspace}{|C-D|}$
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR.
2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutations. Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3: Average Generation Time:  measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
2.5. Mutation Generation
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation. We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
2.5.1. Mutation Generation via Traditional Approaches
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and $\mu$Bert), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that for all mutation generation approaches, we employ the same setting for the code context.
2.5.2. Mutation Generation via LLMs
Recently, LLMs have shown great potential in numerous code-related tasks, to better understand their ability in mutation testing, we also study LLMs for generating mutations.
LLMs operate based on prompts, which consist of natural language instructions, often accompanied by code snippets. LLMs are driven by prompts, which are natural language instructions maybe with code. To effectively leverage LLMs for mutation generation, we design a well-crafted prompt template, shown as Figure 1. We follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
Figure 1. The Prompt Template for LLMs
Table 3. Default Few-Shot Examples from QuixBugs
Correct Version Buggy Version
n = (n & (n - 1)); n = (n ^ (n - 1));
while (!queue.isEmpty()) while (true)
return depth==0; return true;
ArrayList r = new ArrayList();
r.add(first).addll(subset);
to_add(r);
to_add.addAll(subset);
c = bin_op.apply(b,a); c = bin_op.apply(a,b);
while(Math.abs(x-approx*approx)>epsilon) while(Math.abs(x-approx)>epsilon)
3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and $\mu$Bert are over 0.6, and the LLMs from the GPT family are all over 0.7.
Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches. A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
Table 4. Overall Performance of All the Mutation Generation Techniques
Metric PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Mut. # 340728 20831 36505 12946 31656 31945 34417 34741 28058 27614
Mut. Score 0.541 0.521 0.648 0.677 0.702 0.737 0.708 0.714 0.507 0.669
Real Bug Detec. 40.1% 66.9% 63.9% 60.5% 87.0% 89.9% 90.8% 89.4% 38.5% 71.2%
Avg. Ochiai 23.0% 35.5% 32.4% 29.3% 47.7% 56.2% 54.1% 50.9% 26.0% 33.3%
Coupling Rate 23.5% 38.9% 38.1% 47.3% 46.9% 50.6% 49.5% 49.5% 34.0% 45.6%
Comp. Rate 100% 97.6% 44.5% 26.5% 62.5% 76.4% 75.3% 75.8% 16.4% 69.3%
Dup. Mut. Rate 0.0% 0.0% 2.3% 1.7% 10.3% 7.8% 7.0% 7.7% 12.8% 37.1%
A. G. Time (s) 0.02 0.08 3.06 2.34 1.79 1.31 1.65 4.25 7.53 9.06
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in grey background color. GPT-4o-M: GPT-4o-Mini. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat-$\beta$-16b. CL-13b: CodeLlama-Instruct-13b.
3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
3.1.1. RQ1.1 Detectability
As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
Real Bug Real Bug Detection Rate: The row Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches. Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and $\mu$Bert perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively. For LLM-based approaches, the results are significantly more effective than those of traditional ones. GPT-4o-Mini achieves the highest detection rate at 90.8%, closely followed by GPT-4o and DeepSeek, with detection rates of 89.9% and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
Table 5. Number of Real Bugs Detected by Each Mutation Generation Approach
Project Bug # PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Chart 26 19 17 16 10 22 22 23 23 9 8
Lang 65 23 51 27 38 52 55 56 57 11 41
Time 27 12 23 18 15 25 24 26 22 10 17
Math 106 40 82 44 65 92 94 96 95 58 77
Mockito 38 14 28 27 5 30 29 36 34 16 21
Closure 133 38 92 122 103 130 131 131 132 79 114
Cli 39 6 11 27 29 37 39 39 39 2 34
Codec 18 6 11 12 11 14 16 17 17 1 11
Csv 16 7 13 13 7 15 16 16 16 2 13
Gson 18 6 9 11 12 13 16 16 15 1 10
JacksonCore 26 1 17 17 18 23 23 21 21 1 17
Jsoup 93 48 77 90 86 93 92 93 90 57 87
ConDefects 246 121 138 120 116 194 208 203 200 81 156
ALL 851 341 569 544 515 740 765 773 761 328 606
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Particularly, GPT-4o-Mini outperforms other approaches in several projects, including Chart, Time, Mockito, and Jsoup, where it detects the highest number of bugs. Similarly, DeepSeek and GPT-4o demonstrate strong performance, achieving the best results in projects such as Lang, Math, and Closure.
3.1.2. RQ1.2 Coupling
To assess the degree of coupling, we perform comprehensive experiments that include both an overall metric evaluation and a project-wise analysis.
Average Ochiai Coefficient: The row Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and $\mu$BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. The LLM-based approaches demonstrate significantly higher Ochiai coefficients, with GPT-4o achieving the highest value of 56.2%, followed closely by GPT-4o-Mini at 54.1%, and DeepSeek at 50.9%. Overall, GPT-4o and GPT-4o-Mini exhibit their superiority in generating mutations with smaller semantic distances with real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar with Real Bug Detectability, which indicates that the LLM-based appraoches, GPT-4o and GPT-4o-Mini, achieve the highest number of bugs across most projects. DeepSeek also exhibits strong performance, closely following GPT-4o in many cases. All traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and $\mu$BERT across most projects.
Table 6. Number of Bugs with High Ochiai Scores (Over 0.8) Achieved by Each Mutation Generation Approach
Project Bug # PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Chart 26 9 9 3 5 10 13 12 12 7 3
Lang 65 15 34 13 18 32 35 37 34 12 23
Time 27 8 14 8 4 8 12 12 11 4 7
Math 106 20 42 13 27 46 56 48 47 19 33
Mockito 38 1 12 5 3 14 13 17 14 3 3
Closure 133 12 25 40 14 38 69 56 45 21 26
Cli 39 3 3 13 8 18 22 23 20 4 10
Codec 18 3 1 3 3 2 8 6 7 3 2
Csv 16 4 2 1 2 4 10 10 8 3 1
Gson 18 2 1 4 3 2 8 7 6 3 2
JacksonCore 26 2 3 9 8 11 14 12 11 5 6
Jsoup 93 16 6 32 20 46 39 39 41 25 19
ConDefects 246 101 150 132 134 175 179 181 177 112 148
ALL 851 196 302 276 249 406 478 460 433 221 283
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and $\mu$Bert, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, GPT-4o achieves the highest coupling rate, with a value of 50.6%. It is followed closely by GPT-4o-Mini and DeepSeek, with coupling rates of 49.5% and 49.5%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly GPT-4o, in generating highly coupled mutations that better mimic real bugs.
Table 7. Distribution of Coupling Categories Across Mutation Generation Approaches
Category PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Strong 10.1% 9.6% 4.9% 7.8% 7.0% 6.8% 6.2% 6.4% 8.0% 6.8%
Strong Addi. 4.7% 5.8% 4.3% 5.2% 6.7% 10.2% 12.1% 10.3% 5.7% 6.0%
Partial 1.8% 1.1% 0.5% 0.8% 1.2% 1.2% 0.8% 1.0% 2.7% 0.9%
Partial Addi. 6.9% 22.4% 28.4% 33.5% 32.0% 32.5% 30.4% 31.8% 17.6% 31.9%
Not Subs. 30.6% 13.2% 26.7% 20.4% 23.3% 23.1% 21.3% 21.9% 16.7% 21.3%
Not Detected 45.9% 47.9% 35.2% 32.3% 29.8% 26.3% 29.2% 28.6% 49.3% 33.1%
Note: Strong - Strong Substitution, Strong Addi. - Strong Substitution + Additional Failing Tests, Partial - Partial Substitution, Partial Addi. - Partial Substitution + Additional Failing Tests, Not Subs. - Not Substitution.
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as GPT-4o and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. These approaches achieve significantly higher rates in the Partial Substitution + Additional Tests Fail category, with values of 32.5% and 30.4%, respectively. Furthermore, they outperform traditional methods in the Strong Substitution + Additional Tests Fail category, with GPT-4o and GPT-4o-Mini achieving 10.2% and 12.1%, respectively. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
3.1.3. RQ1.3 Diversity
We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
Figure 2. Number of Newly Introduced AST Nodes of the Studied Approaches and the Few-Shot Examples
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang [^3] https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approach, DeepSeek, exhibits the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs are significantly higher than all other approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
Table 8. Ratio of Deletion Mutations and the Top-3 Newly Introduced AST Node Types for Each Approach
Top K Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Del. 15.3% 48.9% 0.4% 5.3% 2.1% 0.5% 1.0% 0.3% 0.1%
1st ST(78.6%) LT(23.2%) LT(29.4%) LT(39.2%) LT(45.4%) LT(43.3%) LT(38.8%) LT(22.7%) LT(47.8%)
2nd LT(21.4%) MR(14.4%) SE(25.7%) BO(17.3%) BO(16.6%) BO(21.6%) BO(19.1%) BO(13.8%) BO(25.7%)
3rd —— CT(9.6%) MR(24.5%) MI(9.9%) CC(8.8%) CC(7.6%) MI(10.1%) MR(13.1%) MI(9.3%)
Del.: The proportion of deletion mutations. LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
Table 9. The Distribution of AST Edit Distance for Each Approach
Distance Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
1 69.2% 58.5% 73.8% 52.4% 50.0% 47.3% 45.1% 36.0% 46.4%
2 27.8% 23.6% 8.9% 36.4% 38.8% 41.9% 44.2% 39.1% 46.6%
3 1.8% 7.3% 2.7% 6.8% 7.2% 6.3% 6.5% 8.0% 3.8%
4 0.6% 3.4% 14.2% 2.4% 2.2% 2.5% 2.6% 6.0% 1.8%
5 0.4% 1.0% 0.1% 0.9% 0.8% 0.8% 0.8% 3.7% 0.5%
$\geq$6 0.2% 6.1% 0.4% 1.0% 1.0% 1.2% 0.8% 7.2% 1.0%
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and $\mu$Bert is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, GPT-4o exhibits a more balanced distribution, with 38.8% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
Answer to RQ1: LLM-based approaches, such as GPT-4o and GPT-4o-Mini, outperform traditional methods in effectiveness. The weighted averages of the rule-based approach in terms of real bug detection rate and coupling rate are 41.64% and 24.38%, while the LLMs are 79.07% ($\uparrow 89.89\%$; 37.43% PPs) and 46.21% ($\uparrow 89.54\%$; 21.83% PPs). Specifically, GPT-4o-Mini achieves the highest real bug detection rate (90.8%) compared to PIT (40.1%) and Major (66.9%). GPT-4o also achieves the highest coupling rate (49.9% vs. 17.9% for PIT and 42.7% for Major). Moreover, LLM-generated mutations demonstrate significantly higher syntactic diversity, with significantly higher new AST node types (e.g., DeepSeek has 49), and a lower proportion of deletions (e.g., 26.3% for GPT-4o vs. 47.9% for Major).
3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., comparability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
3.2.1. RQ2.1: Compilability
The row Comp. Rate of Table 4 presents the overall Compilability Rate (CR) of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% $CR$, and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictedly and previously filters out non-compilable mutations. In contrast, LEAM and $\mu$Bert, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. GPT-4o achieves the highest compilability among LLMs at 76.40%, followed closely by GPT-4o-Mini and DeepSeek, with rates of 75.30% and 75.80%, respectively. Other LLMs such as GPT-3.5 and CodeLlama show moderate $CR$s of 62.50% and 71.20%, respectively, whereas models like StarChat fall behind with a $CR$ of 38.50%. Compared to small models (i.e., LEAM and Bert), GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and $\mu$Bert.
3.2.2. RQ2.2: Duplicate Mutation Rate
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate ($DMR$) is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The row Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% $DMR$. This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and $\mu$Bert, show relatively low $DMR$s of 2.30% and 1.70%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher $DMR$. For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate $DMR$s of 10.30%, 7.80%, and 7.00%, respectively. The open-source LLMs such as StarCoder and DeepSeek demonstrate higher $DMR$s of 12.80% and 7.70%, respectively, while CodeLlama shows the highest $DMR$ at 37.10%.
3.2.3. RQ2.3: Equivalent Mutations
Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity. The procedure for checking a mutation involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent. Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
In total, 959 mutations were labeled, 508 of which were filtered by tests (i.e., killed), and the remaining 451 were manually checked. The entire process took approximately ten hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient ($\kappa$(Viera et al., 2005), achieving an overall $\kappa$ value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
Table 10. Sample Compilable Mutations for Identifying Equivalent Mutations
Metric PIT Major LEAM $\mu$Bert GPT-3.5 GPT-4o GPT-4o-Mini DS-236b SC-16b CL-13b
#Sampled 97 98 96 94 96 96 96 96 94 96
#Test Filtered 26 52 48 53 60 64 59 63 32 51
#Eq. Mut 1 2 2 6 4 3 3 5 10 6
EMR 1.0% 2.0% 2.1% 6.4% 4.2% 3.1% 3.1% 5.2% 10.6% 6.3%
Table 10 presents the estimated Equivalent Mutation Rate ($EMR$), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the $EMR$ for each approach. Among the traditional approaches, PIT achieves the lowest $EMR$ of 1.0%, followed by Major with an $EMR$ of 2.0%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher $EMR$s, with variations across different models. StarChat has the highest $EMR$s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek show relatively lower $EMR$s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini, tend to have slightly higher $EMR$s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated $EMR$ of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutation for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs considering the randomness introduced by sampling.
Answer to RQ2: Rule-based approaches (i.e., PIT and Major) outperform LLMs in comparability rate, duplicate mutation rate, and equivalent mutation rate, with a weighted average of 99.90%, 0.00%, and 1.10% respectively. While the weighted average of LLMs is 63.80% ($\downarrow 36.14\%$; $36.10\%$ PPs), 13.10% ($\uparrow 13.10\%$; $13.10\%$), and 5.30% ($\uparrow 79.25\%$; $4.2\%$ PPs), respectively. Specifically, PIT and Major produce few non-compilable or duplicate mutations, while GPT-4o generates 23.6% non-compilable mutations and 7.8% duplicate mutations. Although rule-based approaches are slightly better in terms of equivalent mutation rates, the equivalent mutation rates across approaches remain comparable, with no significant disparities observed.
3.3. RQ3: Performance on Efficiency
To evaluate the performance on efficiency, we record the average generation time for each approach.
Average Generation Time: The row A. G. Time of Table 4 presents the results of Average Generation Time. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and $\mu$Bert take 3.06 seconds and 2.34 seconds, respectively. Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. While the open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
Answer to RQ3: Rule-based traditional approaches, such as PIT and Major, outperform LLMs in time cost. On average, PIT and Major generate a single mutant in just 0.02s and 0.08s, respectively, whereas the best-performing LLM, GPT-4o, requires 1.31s.
3.4. RQ4: Categorizations of Non-Compilable Mutations
To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
3.4.1. Categorizations of compilation error types
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.
  1. (1)
    Usage of Unknown Methods: this type of error is due to invoking unknown methods under the code context.
  2. (2)
    Code Structural Destruction: this type of error is due to illegal code structures, such as unmatched parentheses or braces.
  3. (3)
    Incorrect Method Parameters: this type of error is due to invoking a method with illegal parameters, including invoking by wrong numbers of parameters or unmatched type.
  4. (4)
    Usage of Unknown Variables: this type of error is due to using unknown variables under the code context.
  5. (5)
    Usage of Unknown Types: this type of error is due to using unknown type names under the code context.
  6. (6)
    Type Mismatch: this type of error is due to using illegal types when applying a certain operation, such as using objects to perform arithmetic operations.
  7. (7)
    Incorrect Initialization: this type of error is due to using illegal forms of constrictors or deleting necessary initialization statements.
  8. (8)
    Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
  9. (9)
    Incorrect Exceptions: this type of error is due to throwing illegal exceptions.
Table 11. Error Types of Non-Compilable Mutations
ID Error Type % ID Error Type %
1 Usage of Unknown Methods 27.3% 6 Type Mismatch 7.6%
2 Code Structural Destruction 22.9% 7 Incorrect Initialization 3.1%
3 Incorrect Method Parameters 12.8% 8 Incorrect Location 3.1%
4 Usage of Unknown Variables 11.2% 9 Incorrect Exceptions 2.3%
5 Usage of Unknown Types 9.6%
Table 11 presents the results. From the table, we observe that the most frequent error type, Usage of Unknown Methods, accounts for 27.3% of total errors, highlighting the challenge of hallucination in LLM-generated code (Huang et al., 2023). The second most common error, Code Structural Destruction, makes up 22.9%, indicating that maintaining syntactic correctness remains a persistent difficulty for LLMs. These findings underscore the need for further improvements in LLMs to enhance their reliability and accuracy in code generation.
(a) Major
(b) LEAM
(c) $\mu$Bert
(d) GPT-3.5
(e) GPT-4o
(f) GPT-4o-M
(g) DS-236b
(h) SC-16b
(i) CL-13b
Figure 3. Distribution of the AST Node Types of Origin Code for Non-Compilable Mutations
3.4.2. Categorizations of code context of non-compilable mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types. For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% ($\mu$Bert), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation. Thus for LLMs, deletion is not the major reason for non-compilation.
Answer to RQ4: The mutations generated by GPTs have nine main types of compilation errors, with Usage of Unknown Methods and Code Structural Destruction being the most prevalent. Deletion mutations are not the primary cause of non-compilable code. An analysis of the surrounding code context where GPTs produce non-compilable mutations reveals that member references and method invocations are the most frequently occurring code features associated with these errors.
4. Discussion
In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
4.1. Sensitivity to Experimental Settings
We discuss the experiment setups that may affect the evaluation results, particularly the setups related to LLMs.
4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutation generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get four types of prompts, shown as follows:
  1. (1)
    Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
  2. (2)
    Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
  3. (3)
    Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
  4. (4)
    Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
Table 12. Comparative Evaluation of Different Prompts for LLMs
Metric GPT-3.5 DS-236b
P1 P2 P3 P4 P1 P2 P3 P4
Mut. # 3681 3688 3350 3210 3803 3845 3802 3804
Mut. Score 0.773 0.785 0.842 0.801 0.784 0.778 0.793 0.814
Real Bug Detec. 79.4% 78.4% 72.2% 72.5% 84.6% 84.0% 81.4% 82.6%
Avg. Ochiai 63.1% 61.8% 60.8% 62.1% 67.3% 64.1% 65.0% 63.7%
Coupling Rate 52.1% 52.2% 47.4% 49.6% 53.1% 51.9% 52.9% 50.7%
Compilability Rate 66.9% 66.8% 60.0% 61.2% 79.1% 78.9% 74.9% 78.2%
Dup. Mut. Rate 11.4% 13.3% 14.9% 12.7% 10.9% 9.4% 9.8% 12.7%
Avg. Gen. Time 1.789 1.782 1.809 1.884 4.252 4.254 4.199 4.397
Table 12 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek across multiple metrics. Among the four prompts, P1 consistently demonstrates superior performance across both models, achieving the highest Compilability Rates (66.9% for GPT-3.5 and 79.1% for DeepSeek) and the best Real Bug Detectability Rates (79.4% and 84.6%, respectively). P1 also yields the highest Avgerage Ochiai Coefficient across both models. These results overall indicate that providing LLMs with well-structured and appropriate information significantly enhances their performance (e.g., P1), whereas excessive information can lead to a decline in effectiveness (e.g., P4).
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 14. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05.
4.1.2. Experiment Setup of the Context Length
Table 13. Performance of Different Context Lengths
Metric 1-Line Context 2-Line Context 3-Line Context
Real Bug Dectc. 85.0% 83.3% 83.3%
Avg. Ochiai 33.3% 35.0% 30.0%
Coupling Rate 36.6% 35.2% 37.1%
Compilability Rate 62.9% 63.2% 67.2%
Dup. Mut. Rate 10.8% 12.5% 11.9%
Table 14. Performance Similarity of Different LLM Configurations
Settings Spea. Coef. $p$-value Pear. Coef. $p$-value
GPT-3.5 P1 v.s. GPT-3.5 P2 0.9762 3.31E-05 0.9999 1.35E-31
GPT-3.5 P1 v.s. GPT-3.5 P3 0.9040 2.00E-03 0.9999 8.66E-23
GPT-3.5 P1 v.s. GPT-3.5 P4 0.9520 2.60E-04 0.9999 1.45E-25
DS-236b P1 v.s. DS-236b P2 1.0000 0.00E+00 0.9999 1.94E-31
DS-236b P1 v.s. DS-236b P3 0.9999 3.31E-05 0.9999 3.71E-30
DS-236b P1 v.s. DS-236b P4 0.9762 3.31E-05 0.9999 1.83E-27
1-Line v.s. 2-Line 0.9762 3.31E-05 0.9980 2.12E-08
2-Line v.s. 3-Line 0.9762 3.31E-05 0.9997 1.04E-10
1-Line v.s. 3-Line 0.9524 2.60E-04 0.9978 2.69E-08
QB-6a (Default) v.s. QB-3 0.9701 6.55E-05 0.9996 1.70E-10
QB-6a (Default) v.s. QB-3b 0.9762 3.31E-05 0.9984 1.00E-08
QB-6a (Default) v.s QB-9 0.9762 3.31E-05 0.9985 7.98E-09
QB-6a (Default) v.s D4J-6 0.9524 2.60E-04 0.9980 2.13E-08
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 13. Similarly, the middle part of Table 14 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95 and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
4.1.3. Impact of Different Few-Shot Examples
Table 15. Impact on Different Few-Shot Examples
Metric QB-6a (Default) QB-3 QB-6b QB-9 D4J-6
Real Bug Det. 0.917 0.933 0.917 0.917 0.900
Avg. Ochiai 0.467 0.450 0.434 0.434 0.367
Coupling Rate 0.136 0.128 0.132 0.121 0.139
Comp. Rate 65.7% 62.7% 63.9% 61.8% 62.1%
Dup Mut. Rate 13.7% 12.8% 12.8% 13.8% 12.5%
Equivalent Mutation Rate 1.3% 1.6% 1.3% 1.1% 1.6%
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 15.
We further compute their similarity, shown as the bottom part of Table 14. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude that although different few-shot examples have perturbation, their performance is very similar.
4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 11397. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 16. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
Table 16. Performance Under the Same Number of Mutations
Metric PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Real Bug Det. 31.6% 68.8% 64.6% 66.0% 82.2% 87.4% 88.6% 87.0% 33.9% 68.5%
Coupling Rate 17.4% 45.4% 37.2% 43.6% 45.6% 48.8% 47.9% 47.9% 27.1% 43.6%
Ochiai Coefficient 12.6% 22.2% 14.1% 19.0% 26.5% 36.0% 31.8% 34.1% 15.4% 16.9%
Comp. Rate 96.9% 54.1% 26.1% 61.2% 75.5% 74.3% 74.6% 12.4% 69.5%
Dup Mut. Rate 0.0% 2.6% 1.9% 9.9% 7.7% 7.1% 7.6% 9.6% 38.3%
4.2. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches, and suggest potential improvements.
4.2.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations. For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations. In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
// Mutation-1
-   Size2D size = this.rightBlock.arrange(g2, c4);
+   Size2D size = this.leftBlock.arrange(g2, c4);
// Mutation-2
-   else {
-     g2.setStroke(getItemOutlineStroke(row, column, selected));
+   else if (seriesCount == 1) {
+     g2.setStroke(getItemOutlineStroke(row, column, selected));
Listing 1: Two Example Mutations of GPT-3.5 on Chart-1
Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
4.2.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
4.2.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore integrating LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
5. Threats to Validity
In this section, we illustrate the threats to the validity of our study.
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek (66), StarChat (13), CodeLlama (28), LEAM (287), $\mu$Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), $\mu$Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs,
The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
6. Related Work
In this section, we introduce the related work of our study.
6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., $a+b\mapsto a-b$). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed $\mu$Bert (Degiovanni and Papadakis, 2022). $\mu$Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al.(Dakhel et al., 2024) and Ibrahimzada et al.(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.
6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo, GPT-4o, and GPT-4o-Mini are closed-source LLMs released by OpenAI (Achiam et al., 2023), while DeepSeek-Coder, StarChat, and CodeLlama are open-source. An open-source LLM is usually fine-tuned against a base model for specified tasks. DeepSeek-Coder (Zhu et al., 2024) is based on DeepSeek. CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets. StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.