### v4

# A Comprehensive Study on Large Language Models for Mutation Testing

Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we conduct a comprehensive empirical study evaluating BugFarm and LLMorpheus (the two state-of-the-art LLM-based approaches), alongside seven LLMs using our newly designed prompt, including both leading open- and closed-source models, on 851 real bugs from two Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection. That is, 87.98% (for LLMs) vs. 41.64% (for rule-based); an increase of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In the most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Harman et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021b; Petrović and Ivanković, 2018) reported on results from the deployment of mutation testing at their companies. In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique, is challenging because of the following three important issues:
Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed. Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper examines how well Large-Language Models (LLMs) address these challenges, in comparison to more traditional rule-based approaches. We conducted a comprehensive empirical study which studies BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutation generation approaches), alongside seven LLMs adopting our newly designed prompt, covering both leading open- and closed-source models, and four existing mutation testing techniques (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)) with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our primary experiments required the generation of 666,148 mutations.
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies, turbulence on prompt, and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), DeepSeek-V3 (Liu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. We also explore the potential of using LLMs for generating high-order mutations. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This results in a higher fault detection rate compared to traditional approaches. For instance, the mutants generated by DeepSeek-V3-761b and GPT-4o-Mini achieve fault detection rates of 91.1 % and 90.8%, respectively, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that the DeepSeek models exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The existing LLM-based approach, BugFarm, introduces 31 node types, higher than the traditional approaches as well. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also show that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have an AST edit distance of one, whereas for GPT-4o, only 50% of its mutations have the same distance.
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
We analyze the alive (i.e., survive) mutations of each approach, which are important to strengthen the test case. For one alive non-equivalent mutation, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection. We analyze the distribution of alive mutations across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutations, which consequently remain alive, providing more valuable guidance for enhancing unit tests.
Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
In summary, our paper makes the following main contributions:
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our study not only covers other existing LLM-based mutation generation approaches, but also introduces a new effective prompt template. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
We construct a high-quality dataset of Java mutations, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
The remainder of this paper is organized as follows. Section 2 presents the design of our study, including the research questions, studied approaches, datasets, and metrics. Section 3 presents the empirical evaluation of these generation approaches and answers to each research question. Section 4 discusses the influences of different experimental setups and the implications summarized from the findings. Section 5 presents the threats to the validity of our study. Section 6 discusses the related work. Section 7 concludes this work.
2. Study Design
In this section, we introduce how our study is designed and conducted.
2.1. Overview and Research Questions
This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutations in representing real bugs. Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of mutations generated by LLMs compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
RQ3 (Efficiency) What are the general costs of different mutation generation approaches? This research question investigates the general costs of different mutation generation approaches. We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutation generation approaches.
RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches. Understanding these errors can guide mutation generation approaches to enhance their practical usability.
RQ5 (Alive Mutations) What are the distributions of alive mutations across different approaches? This RQ aims to investigate the underlying reasons why mutations remain alive, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests. By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutations and their potential value for guiding test suite improvement.
2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches are for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. The details of the settings for each generation approach are illustrated in Section 2.5. Table 1 summarizes the mutation generation approaches covered by this study.
| Type | Name | Description |
| Traditional | PIT | Widely studied mutation testing framework with 29 mutation operators. |
| Major | Widely studied mutation testing framework with 11 mutation operators. | |
| LEAM | Custom-trained models learning from a large corpus of real bugs and patches. | |
| BERT | Mutation generation with the pre-trained model CodeBERT. | |
| LLM-based | BF (GPT-4o) | The existing LLM-based approach BugFarm with GPT-4o. |
| LLMph (GPT-4o) | The existing LLM-based mutation generation approach LLMorpheus with GPT-4o. | |
| LLMut (GPT-3.5) | Our prompt with GPT-3.5-Turbo, trained on data before 2021/09, released in 2021/03. | |
| LLMut (GPT-4o) | Our prompt with GPT-4o, trained on data before 2023/10, released in 2024/05. | |
| LLMut (GPT-4o-M) | Our prompt with GPT-4o-Mini, trained on data before 2023/10, released in 2024/07. | |
| LLMut (SC-16b) | Our prompt with StarChat--16b, whose base model is StarCoder, released in 2023/06. | |
| LLMut (CL-13b) | Our prompt with CodeLlama-Instruct-13b, whose base model is Llama, released in 2023/08. | |
| LLMut (DS-236b) | Our prompt with DeepSeek-Coder-V2-236b, whose base model is DeepSeek, released in 2024/07. | |
| LLMut (DS-671b) | Our approach with DeepSeek-V3-761b, whose base model is DeepSeek, released in 2024/12. |
2.2.1. Traditional Approaches
For Java, there exist several mutation testing frameworks, such as JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), PIT (Coles et al., 2016), and Major (Just et al., 2011). As some tools are no longer maintained, following existing studies (Kaufman et al., 2022; Ojdanic et al., 2023; Kushigian et al., 2024), we employ PIT (Coles et al., 2016), Major (Just et al., 2011), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022). Major and PIT are widely studied rule-based approaches, while LEAM and BERT involve deep learning models.
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022, [n. d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
BERT: BERT (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, BERT (Devlin et al., 2018). Equipped with the pre-trained model, BERT transforms the task of mutation generation to code completion under the surrounding context.
2.2.2. LLM-based Approaches
Our study incorporates two existing LLM-based mutation generation approaches (i.e., BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutation generation, which will be presented later.
Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutation generation approaches, BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025). BugFarm trains a small-scale model for selecting lines for mutation, and then utilizes LLMs to mutate the target line. LLMorpheus performs mutation given the whole method as context. Note that LLMorpheus was originally designed for JavaScript. Upon examining its prompt, we found it to be language-agnostic. Therefore, we reimplemented the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutation) while retaining the original prompt. For both approaches, we employ GPT-4o (Achiam et al., 2023), the current most widely used model, as their backend LLM.
LLMs with Our Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutation generation, shown as Figure 1. For ease of presentation, in the following part of this paper, we call our prompt LLMut.
To craft effective prompts, we follow the best practices, which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. The placeholder MUT_NUMBER specifies the number of mutations required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated. In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs. We further explore the different prompt settings in Section 4.1 and the impact of prompt turbulence in Section 4.2.
To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 2. As LLM capabilities continue to advance, the optimal number of few-shot examples may vary. We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3.
| Correct Version | Buggy Version |
| n = (n & (n - 1)); | n = (n ˆ (n - 1)); |
| while (!queue.isEmpty()) | while (true) |
| return depth==0; | return true; |
ArrayList r = new ArrayList();
r.add(first).addll(subset); to_add(r); |
to_add.addAll(subset); |
| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
| while(Math.abs(x-approx*approx)>epsilon) | while(Math.abs(x-approx)>epsilon) |
To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
Compared with the prompts used in BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm and LLMorpheus333https://github.com/githubnext/llmorpheus, our prompt design offers several advantages:
it incorporates real-world bugs as few-shot examples, which have been shown to significantly enhance LLM performance in many code-related tasks;
it determines the number of generated mutants based on the size of the target code, rather than on a fixed number;
it specifies the mutation location using the target statement itself rather than line numbers, enabling more precise localization (Fan et al., 2023a); and
it employs JSON output formatting, which is widely used in LLM fine-tuning and downstream tasks.
Based on LLMut, our study comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), DS-236b (DeepSeek-Coder-V2-236b), and DS-671b (DeepSeek-V3-671b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 671 billion. Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora. This selection allows us to evaluate mutation generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
2.3. Datasets
| Dataset | Project | # of Bugs | Time Span |
| Defects4J (D4J) | Math | 106 | 2006/06/05 - 2013/08/31 |
| Lang | 65 | 2006/07/16 - 2013/07/07 | |
| Chart | 26 | 2007/07/06 - 2010/02/09 | |
| Time | 27 | 2010/10/27 - 2013/12/02 | |
| Closure | 133 | 2009/11/12 - 2013/10/23 | |
| Mockito | 38 | 2009/06/20 - 2015/05/20 | |
| Cli | 39 | 2007/05/15 - 2018/02/26 | |
| Codec | 18 | 2008/04/27 - 2017/03/26 | |
| Csv | 16 | 2012/03/27 - 2018/05/18 | |
| Gson | 18 | 2010/11/02 - 2017/09/21 | |
| JacksonCore | 26 | 2013/08/28 - 2019/04/05 | |
| Jsoup | 93 | 2011/07/02 - 2019/07/04 | |
| ConDefects (CD) | — | 246 | 2024/03/01 - 2024/06/30 |
| Total | — | 851 | 2006/07/16 - 2024/06/30 |
As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 3. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014c; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains historical bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from Defects4J 2.0, which contain 605 bugs in total. However, from Table 1 and Table 3, we observe that the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder444https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs. Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it is released after the ConDefects dataset.
2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives, as mentioned before. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus, we denote the compilable mutations as the set . Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ). Therefore, we have . The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contribution to discovering the weakness of tests. However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025). Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ). The mutation score is then computed as follows:
Note that traditional mutation generation approaches, such as Major and PIT, produce very few duplicate mutations (i.e., is 0). Hence, early studies (Kintis et al., 2017) defined equivalent mutations as those semantically equivalent to the original code, which may implicitly include , which is an empty set. However, LLM-based approaches tend to generate a large number of non-mutated duplicates. Moreover, while duplicate mutations can be easily filtered, identifying equivalent mutations remains challenging (Tian et al., 2024). Therefore, in this paper, we refine the definition of equivalent mutations to be those that are syntactically different but semantically equivalent to the original code.
2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed below.
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014c), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al. (Ojdanic et al., 2023).
More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug . Given a bug , its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., ), shown as the following formula:
Given a mutation generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutations are considered to have high Ochiai if they satisfy . This threshold indicates a strong semantic similarity between the generated mutations and the real bug. For each mutation generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ). Note that Real Bug Detection Rate and Coupling Rate measure effectiveness from different perspectives. Real Bug Detection Rate evaluates effectiveness at the bug level, i.e., the proportion of bugs that can be revealed by at least one mutant. In contrast, Coupling Rate evaluates effectiveness at the mutation level, i.e., the proportion of mutants whose failing tests overlap with bug-triggering tests. For instance, a mutation approach may generate a large number of high-quality mutants for only a few specific bugs, but few mutants for others. This can result in a high Coupling Rate (since the generated mutants are strongly aligned with the bug-triggering tests of those few bugs) but a low Real Bug Detection Rate (since many other bugs remain undetected). Conversely, another approach may generate mutants that cover a wide range of bugs, thereby achieving a high Real Bug Detection Rate. However, if the mutants for each bug exhibit only weak coupling with the bug-triggering tests, the overall Coupling Rate will remain low. Therefore, these two metrics are complementary, with each reflecting only one aspect of mutation effectiveness.
RQ1.2: Coupling Categorization: We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutation (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
Partial Substitution: Some triggering tests fail and no additional tests fail.
Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
Not Substitution: Only non-triggering tests fail.
Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
Later, we will classify the mutations of each approach and analyze their distributions.
RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. In other words, we identify the unique AST node types introduced by each approach’s generated mutants, relative to the set of node types present in the original program. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes. The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics. However, it still has inherent limitations, such as its inability to fully reflect the impact of mutations on program behavior. We further discuss these limitations later in Section 5.
2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, which is computed as:
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher compilability rate.
RQ2.2: Duplicate Mutation Rate: the proportion of duplicate mutations, computed as:
LLMs sometimes generate duplicate mutations, i.e., mutations that are either identical to the original code or identical to each other. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data. After eliminating duplicate mutations, all remaining mutations are guaranteed to be syntactically distinct both from the original code and from each other.
RQ2.3: Equivalent Mutation Rate: the proportion of equivalent mutation, computed as:
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR. Note that while equivalent mutations are also considered semantically duplicate, they are not easily identifiable. Therefore, we distinguish them from other duplicate mutations, which can be filtered syntactically.
2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutations. Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3.1: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
RQ3.2: Average Token: measures the average number of consumed tokens per mutation, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time). Since input and output token costs differ across models, we calculate them separately.
2.5. Mutation Generation Setups
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation.
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that we supply the same code snippets as context across all mutation generation approaches to ensure a fair comparison.
For all LLM-based approaches, we use the default settings of each model (e.g., token limits and temperature). For models accessible via Web API (e.g., GPT series and DeepSeek models), we rely on their official APIs. For the remaining open-source LLMs, we perform inference on four rented cloud servers, each equipped with dual NVIDIA GeForce RTX 3090 Ti GPUs, 168 GB of memory, and a 64-core Intel(R) Xeon(R) Platinum 8350C processor.
3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and BERT are over 0.6, and the LLMs from the GPT family are all over 0.7.
Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches. A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
| Name | #Mut. | Mut. S. | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate | Time |
| PIT | 340728 | 0.541 | 40.1% | 23.0% | 23.5% | 100% | 0.0% | 0.02 |
| Major | 20831 | 0.521 | 66.9% | 35.5% | 38.9% | 97% | 0.0% | 0.08 |
| LEAM | 36505 | 0.648 | 63.9% | 32.4% | 38.1% | 44.5% | 2.3% | 3.06 |
| BERT | 12946 | 0.677 | 60.5% | 29.3% | 47.3% | 26.5% | 1.7% | 2.34 |
| BF (GPT-4o) | 13068 | 0.762 | 58.6% | 17.7% | 47.9% | 74.1% | 7.5% | 5.35 |
| LLMph (GPT-4o) | 20043 | 0.754 | 68.5% | 24.1% | 51.7% | 67.8% | 7.1% | 3.41 |
| LLMut (GPT-3.5) | 31656 | 0.702 | 87.0% | 47.7% | 46.9% | 62.5% | 10.3% | 1.79 |
| LLMut (GPT-4o) | 31945 | 0.737 | 89.9% | 56.2% | 50.6% | 76.4% | 7.8% | 1.31 |
| LLMut (GPT-4o-M) | 34417 | 0.708 | 90.8% | 54.1% | 49.5% | 75.3% | 7.0% | 1.65 |
| LLMut (DS-761b) | 33596 | 0.749 | 91.1% | 59.8% | 50.8% | 77.6% | 7.5% | 2.57 |
| LLMut (DS-236b) | 34741 | 0.714 | 89.4% | 50.9% | 49.5% | 75.8% | 7.7% | 4.25 |
| LLMut (SC-16b) | 28058 | 0.507 | 38.5% | 21.7% | 34.0% | 16.4% | 12.8% | 7.53 |
| LLMut (CL-13b) | 27614 | 0.669 | 71.2% | 33.3% | 45.6% | 69.3% | 37.1% | 9.06 |
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in a grey background color.
BF: BugFarm. LLMph: LLMorpheus. LLMut: Our new prompt.
GPT-4o-M: GPT-4o-Mini. DS-761b: DeepSeek-V3-761b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
3.1.1. RQ1.1 Detectability
As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and BERT perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively.
For LLM-based approaches, the results are significantly more effective than those of traditional ones. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve a rate of 58.6% and 68.5%, respectively. For the LLMs based on our prompting, DeepSeek-V3-761b achieves the highest detection rate at 91.1%, closely followed by GPT-4o-Mini, GPT-4o and DeepSeek-Coder-V2-236b, with detection rates of 90.8%, 89.9%, and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
| PIT | 19 | 23 | 12 | 40 | 14 | 38 | 6 | 6 | 7 | 6 | 1 | 48 | 121 | 341 |
| Major | 17 | 51 | 23 | 82 | 28 | 92 | 11 | 11 | 13 | 9 | 17 | 77 | 138 | 569 |
| LEAM | 16 | 27 | 18 | 44 | 27 | 122 | 27 | 12 | 13 | 11 | 17 | 90 | 120 | 544 |
| BERT | 10 | 38 | 15 | 65 | 5 | 103 | 29 | 11 | 7 | 12 | 18 | 86 | 116 | 515 |
| BF (GPT-4o) | 14 | 37 | 15 | 57 | 28 | 93 | 23 | 6 | 8 | 11 | 7 | 83 | 117 | 499 |
| LLMph (GPT-4o) | 17 | 51 | 20 | 61 | 31 | 121 | 30 | 12 | 8 | 12 | 9 | 87 | 124 | 583 |
| LLMut (GPT-3.5) | 22 | 52 | 25 | 92 | 30 | 130 | 37 | 14 | 15 | 13 | 23 | 93 | 194 | 740 |
| LLMut (GPT-4o) | 22 | 55 | 24 | 94 | 29 | 131 | 39 | 16 | 16 | 16 | 23 | 92 | 208 | 765 |
| LLMut (GPT-4o-M) | 23 | 56 | 26 | 96 | 36 | 131 | 39 | 17 | 16 | 16 | 21 | 93 | 203 | 773 |
| LLMut (DS-761b) | 24 | 61 | 27 | 101 | 35 | 131 | 39 | 13 | 16 | 16 | 25 | 92 | 195 | 775 |
| LLMut (DS-236b) | 23 | 57 | 22 | 95 | 34 | 132 | 39 | 17 | 16 | 15 | 21 | 90 | 200 | 761 |
| LLMut (SC-16b) | 9 | 11 | 10 | 58 | 16 | 79 | 2 | 1 | 2 | 1 | 1 | 57 | 81 | 328 |
| LLMut (CL-13b) | 8 | 41 | 17 | 77 | 21 | 114 | 34 | 11 | 13 | 10 | 17 | 87 | 156 | 606 |
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Notably, DeepSeek-V3-761b detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count. Similarly, DeepSeek-Coder-V2-236b, GPT-4o, and GPT-4o-Mini demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
3.1.2. RQ1.2 Coupling
To assess the degree of coupling, we perform comprehensive experiments that include both an overall metric evaluation and a project-wise analysis.
Average Ochiai Coefficient: The column Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. Existing other LLM-based approaches achieve an overall lower Ochiai coefficient of 17.7% and 24.1%, respectively. The LLM-based approaches using our prompt achieve notably higher Ochiai coefficients. DeepSeek-V3-761b attains the highest value at 59.8%, followed by GPT-4o at 56.2% and GPT-4o-Mini at 54.1%. The LLMs based on our prompt outperform other existing methods, not only traditional approaches but also existing LLM-based ones, showing their superiority in generating mutations that are semantically closer to real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar to Real Bug Detectability, which indicates that the LLM DeepSeek-V3-761b, based on our prompt, achieves the highest number of bugs across most projects. The closed-source LLMs GPT-4o and GPT-4o-Mini also exhibit strong performance, closely following DeepSeek-V3-761b in many cases. Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
| PIT | 9 | 15 | 8 | 20 | 1 | 12 | 3 | 3 | 4 | 2 | 2 | 16 | 101 | 196 |
| Major | 9 | 34 | 14 | 42 | 12 | 25 | 3 | 1 | 2 | 1 | 3 | 6 | 150 | 302 |
| LEAM | 3 | 13 | 8 | 13 | 5 | 40 | 13 | 3 | 1 | 4 | 9 | 32 | 132 | 276 |
| BERT | 5 | 18 | 4 | 27 | 3 | 14 | 8 | 3 | 2 | 3 | 8 | 20 | 134 | 249 |
| BF (GPT-4o) | 9 | 15 | 3 | 13 | 1 | 13 | 3 | 2 | 1 | 4 | 2 | 14 | 71 | 151 |
| LLMph (GPT-4o) | 6 | 25 | 7 | 24 | 6 | 22 | 9 | 2 | 3 | 3 | 5 | 12 | 81 | 205 |
| LLMut (GPT-3.5) | 10 | 32 | 8 | 46 | 14 | 38 | 18 | 2 | 4 | 2 | 11 | 46 | 175 | 406 |
| LLMut (GPT-4o) | 13 | 35 | 12 | 56 | 13 | 69 | 22 | 8 | 10 | 8 | 14 | 39 | 179 | 478 |
| LLMut (GPT-4o-M) | 12 | 37 | 12 | 48 | 17 | 56 | 23 | 6 | 10 | 7 | 12 | 39 | 181 | 460 |
| LLMut (DS-761b) | 13 | 40 | 15 | 62 | 21 | 79 | 27 | 8 | 10 | 12 | 17 | 59 | 146 | 509 |
| LLMut (DS-236b) | 12 | 34 | 11 | 47 | 14 | 45 | 20 | 7 | 8 | 6 | 11 | 41 | 177 | 433 |
| LLMut (SC-16b) | 7 | 12 | 4 | 19 | 3 | 21 | 4 | 3 | 3 | 3 | 5 | 25 | 112 | 221 |
| LLMut (CL-13b) | 3 | 23 | 7 | 33 | 3 | 26 | 10 | 2 | 1 | 2 | 6 | 19 | 148 | 283 |
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, LLMorpheus achieves the highest coupling rate, with a value of 51.7%. It is followed closely by DeepSeek-V3-761b and GPT-4o-Mini based on our prompt, with coupling rates of 50.8% and 50.6%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus, DeepSeek-V3-761b, and GPT-4o, in generating overall highly coupled mutations that better mimic real bugs.
| Category | Strong | Strong Addi | Partial | Partial Addi | Not Subs. | Not Detected |
| PIT | 10.1% | 4.7% | 1.8% | 6.9% | 30.6% | 45.9% |
| Major | 9.6% | 5.8% | 1.1% | 22.4% | 13.2% | 47.9% |
| LEAM | 4.9% | 4.3% | 0.5% | 28.4% | 26.7% | 35.2% |
| BERT | 7.8% | 5.2% | 0.8% | 33.5% | 20.4% | 32.3% |
| BugFarm (GPT-4o) | 8.4% | 5.4% | 0.6% | 33.6% | 28.3% | 23.8% |
| LLMorpheus (GPT-4o) | 9.3% | 4.9% | 2.1% | 35.4% | 23.6% | 24.6% |
| LLMut (GPT-3.5) | 7.0% | 6.7% | 1.2% | 32.0% | 23.3% | 29.8% |
| LLMut (GPT-4o) | 6.8% | 10.2% | 1.2% | 32.5% | 23.1% | 26.3% |
| LLMut (GPT-4o-M) | 6.2% | 12.1% | 0.8% | 30.4% | 21.3% | 29.2% |
| LLMut (DS-761b) | 11.2% | 4.6% | 1.9% | 33.1% | 24.1% | 25.1% |
| LLMut (DS-236b) | 6.4% | 10.3% | 1.0% | 31.8% | 21.9% | 28.6% |
| LLMut (SC-16b) | 8.0% | 5.7% | 2.7% | 17.6% | 16.7% | 49.3% |
| LLMut (CL-13b) | 6.8% | 6.0% | 0.9% | 31.9% | 21.3% | 33.1% |
Note: Strong - Strong Substitution, Strong Addi. - Strong Substitution + Additional Failing Tests, Partial - Partial Substitution, Partial Addi. - Partial Substitution + Additional Failing Tests, Not Subs. - Not Substitution.
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate, indicating these mutations semantically overlap with their corresponding real bugs. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as DeepSeek-V3-761b, GPT-4o, and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. For instance, DeepSeek-V3-761b achieves the highest proportion in the Strong category at 11.2%, while GPT-4o-Mini leads in the Strong + Additional Tests Fail category, with 12.1%. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
3.1.3. RQ1.3 Diversity
We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approaches, DeepSeek-V3-761b and DeepSeek-Coder-V2-236b, exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs based on our prompt are significantly higher than all other approaches in AST node types. The existing LLM-based approaches, BugFarm and LLMorpheus, introduce 31 and 21 new AST node types, respectively. While BugFarm outperforms traditional approaches, it still lags behind ours. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
| Approach | Deletion | 1st | 2nd | 3rd |
| Major | 15.3% | ST (78.6%) | LT (21.4%) | —— |
| LEAM | 48.9% | LT (23.2%) | MR (14.4%) | CT (9.6%) |
| BERT | 0.4% | LT (29.4%) | SE (25.7%) | MR (24.5%) |
| BugFarm (GPT-4o) | 5.5% | LT (41.9%) | BO (12.2%) | MI (9.0%) |
| LLMorpheus (GPT-4o) | 0.3% | LT (24.0%) | BO (23.7%) | MR (15.8%) |
| LLMut (GPT-3.5) | 5.3% | LT (39.2%) | BO (17.3%) | MI (9.9%) |
| LLMut (GPT-4o) | 2.1% | LT (45.4%) | BO (16.6%) | CC (8.8%) |
| LLMut (GPT-4o-M) | 0.5% | LT (43.3%) | BO (21.6%) | CC (7.6%) |
| LLMut (DS-761b) | 3.8% | LT (44.2%) | BO (15.6%) | CC (8.6%) |
| LLMut (DS-236b) | 1.0% | LT (38.8%) | BO (19.1%) | MI (10.1%) |
| LLMut (SC-16b) | 0.3% | LT (22.7%) | BO (13.8%) | MR (13.1%) |
| LLMut (CL-13b) | 0.1% | LT (47.8%) | BO (25.7%) | MI (9.3%) |
Deletion: The proportion of deletion mutations. BO: BinaryOperation. CC: ClassCreator. CT: Cast. LT: Literal. MI: MethodInvocation. MR: MemberReference. SE: StatementExpression. ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutations deleting the target code element and the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., ). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
| Distance | 1 | 2 | 3 | 4 | 5 | 6 |
| Major | 69.2% | 27.8% | 1.8% | 0.6% | 0.4% | 0.2% |
| LEAM | 58.5% | 23.6% | 7.3% | 3.4% | 1.0% | 6.1% |
| BERT | 73.8% | 8.9% | 2.7% | 14.2% | 0.1% | 0.4% |
| BugFarm (GPT-4o) | 44.7% | 43.5% | 8.3% | 1.6% | 0.6% | 1.4% |
| LLMorpheus (GPT-4o) | 54.2% | 25.7% | 14.6% | 3.7% | 0.9% | 0.8% |
| LLMut (GPT-3.5) | 52.4% | 36.4% | 6.8% | 2.4% | 0.9% | 1.0% |
| LLMut (GPT-4o) | 50.0% | 38.8% | 7.2% | 2.2% | 0.8% | 1.0% |
| LLMut (GPT-4o-M) | 47.3% | 41.9% | 6.3% | 2.5% | 0.8% | 1.2% |
| LLMut (DS-761b) | 46.1% | 44.2% | 6.8% | 1.5% | 0.4% | 0.9% |
| LLMut (DS-236b) | 45.1% | 44.2% | 6.5% | 2.6% | 0.8% | 0.8% |
| LLMut (SC-16b) | 36.0% | 39.1% | 8.0% | 6.0% | 3.7% | 7.2% |
| LLMut (CL-13b) | 46.4% | 46.6% | 3.8% | 1.8% | 0.5% | 1.0% |
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, DeepSeek-V3-761b exhibits a more balanced distribution, with 44.2% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
3.2.1. RQ2.1: Compilability
The column Comp. Rate of Table 4 presents the overall Compilability Rate () of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictively and previously filters out non-compilable mutations. In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve relatively high rates of 74.1% and 67.8%, respectively. Among the LLMs using our prompt, DeepSeek-V3-761b attains the highest at 77.6%, followed closely by GPT-4o (76.4%), DeepSeek-Coder-V2-236b (75.8%), and GPT-4o-Mini (75.3%). Compared to small models (i.e., LEAM and BERT), GPTs and DeepSeek models show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and BERT.
3.2.2. RQ2.2: Duplicate Mutation Rate
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The column Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher . For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.3%, 7.8%, and 7.0%, respectively. For the existing LLM-based approaches, BugFarm and LLMorpheus, exhibit similar s of 7.5% and 7.1%, respectively. The open-source LLMs such as StarCoder and DeepSeek-Coder-V2-236b demonstrate higher s of 12.8% and 7.7%, respectively, while CodeLlama shows the highest at 37.1%.
3.2.3. RQ2.3: Equivalent Mutations
Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity. The procedure for checking a mutation involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent. Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
In total, 1247 mutations were labeled, 709 of which were filtered by tests (i.e., killed), and the remaining 538 were manually checked. The entire process took approximately twelve hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
| Approach | #Sampled | #Test Filtered | #Eq. Mut. | EMR |
| PIT | 97 | 26 | 1 | 1.0% |
| Major | 96 | 52 | 2 | 2.1% |
| LEAM | 96 | 48 | 2 | 2.1% |
| BERT | 94 | 64 | 3 | 6.4% |
| BugFarm (GPT-4o) | 96 | 67 | 3 | 3.1% |
| LLMorpheus (GPT-4o) | 96 | 69 | 4 | 4.2% |
| LLMut (GPT-3.5) | 96 | 60 | 4 | 4.2% |
| LLMut (GPT-4o) | 96 | 64 | 3 | 3.1% |
| LLMut (GPT-4o-M) | 96 | 59 | 3 | 3.1% |
| LLMut (DS-761b) | 96 | 65 | 6 | 6.3% |
| LLMut (DS-236b) | 96 | 63 | 5 | 5.2% |
| LLMut (SC-16b) | 94 | 32 | 10 | 10.6% |
| LLMut (CL-13b) | 96 | 51 | 6 | 6.3% |
Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek-Coder-V2-236b show relatively lower s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutations for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling. Moreover, the definition of equivalent mutations in this paper is the set of syntactically different but semantically equivalent mutations, which may differ from existing studies (Tip et al., 2024).
3.3. RQ3: Performance on Efficiency
To evaluate the performance on efficiency, we record the average generation time for each approach.
Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively. The existing LLM-based approaches, Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. The open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches. Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution. Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations. We further discuss these potential impacts in the Section 5.
| Approach | Avg. Input Token | Avg. Output Token |
| BugFarm (GPT-4o) | 200 | 368 |
| LLMorpheus (GPT-4o) | 370 | 74 |
| LLMut (GPT-3.5) | 126 | 59 |
| LLMut (GPT-4o) | 126 | 59 |
| LLMut (GPT-4o-M) | 126 | 59 |
| LLMut (DS-761b) | 126 | 58 |
| LLMut (DS-236b) | 126 | 59 |
| LLMut (SC-16b) | 126 | 63 |
| LLMut (CL-13b) | 126 | 58 |
Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate one mutation. The existing LLM-based approaches, BugFarm and LLMorpheus, consume 200 and 370 input tokens, and 368 and 74 output tokens per mutation, respectively. LLMs using our prompt consume substantially fewer tokens, averaging 126 input tokens and 59–63 output tokens per mutation. As all LLMs share identical inputs, their input token counts remain the same.
3.4. RQ4: Categorizations of Non-Compilable Mutations
To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
3.4.1. Categorizations of compilation error types
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.
Usage of Unknown Methods: This type of error is due to invoking unknown methods under the code context.
Code Structural Destruction: This type of error is due to illegal code structures, such as unmatched parentheses or braces.
Incorrect Method Parameters: This type of error is due to invoking a method with illegal parameters, including invoking with the wrong number of parameters or mismatched types.
Usage of Unknown Variables: This type of error is due to using unknown variables in the code context.
Usage of Unknown Types: This type of error is due to using unknown type names in the code context.
Type Mismatch: This type of error is due to using illegal types when applying a certain operation, such as using objects to perform arithmetic operations.
Incorrect Initialization: This type of error is due to using illegal forms of constructors or deleting necessary initialization statements.
Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
Incorrect Exceptions: This type of error is due to throwing illegal exceptions.
| ID | Error Type | % | ID | Error Type | % |
| 1 | Usage of Unknown Methods | 27.3% | 6 | Type Mismatch | 7.6% |
| 2 | Code Structural Destruction | 22.9% | 7 | Incorrect Initialization | 3.1% |
| 3 | Incorrect Method Parameters | 12.8% | 8 | Incorrect Location | 3.1% |
| 4 | Usage of Unknown Variables | 11.2% | 9 | Incorrect Exceptions | 2.3% |
| 5 | Usage of Unknown Types | 9.6% | — | — | — |
Table 12 presents the results. From the table, we observe that the most frequent error type, Usage of Unknown Methods, accounts for 27.3% of total errors, highlighting the challenge of hallucination in LLM-generated code (Huang et al., 2023). The second most common error, Code Structural Destruction, makes up 22.9%, indicating that maintaining syntactic correctness remains a persistent difficulty for LLMs. These findings underscore the need for further improvements in LLMs to enhance their reliability and accuracy in code generation.
3.4.2. Categorizations of code context of non-compilable mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types. For example, for GPT and DeepSeek models, about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-761b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation.
3.5. RQ5: Distributions of Alive Mutations
As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than LLMs, indicating their mutations are harder to detect. Therefore, we intend to analyze the reasons for survival mutations. To address this RQ, we adopt a setting similar to the analysis of equivalent mutations. Specifically, for each approach, we randomly sample its alive mutations and manually classify them. We employ the categories of alive mutation causes defined in the study by Just et al. (Just et al., 2014a) as follows.
(1) Not covered, i.e., the mutation is located in code never executed by the test suite. We refer to the mutation set of this category as .
(2) Covered but no state infection, i.e., the mutation is executed but does not propagate any state change. We refer to the mutation set of this category as .
(3) State infection without detection, i.e., the mutation alters the program state but any test oracle does not observe the deviation. We refer to the mutation set of this category as .
These categories are mutually exclusive, with the following relationships: , , and . This classification enables us to analyze the underlying reasons for alive mutations and further evaluate the extent to which different approaches contribute to strengthening the test suite. Given an alive mutation, we first determine whether it is equivalent; if so, it is discarded. For non-equivalent mutations, we then check whether they are uncovered. If a mutation is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified. Finally, the remaining alive mutations that are covered but do not lead to detectable differences are classified as not detected. This stepwise process ensures that every alive mutation is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
| Approach | #Sampled | #Eq. Mut. | NC | NI | ND | NC% | NI% | ND% |
| PIT | 96 | 3 | 43 | 27 | 23 | 46.24% | 29.03% | 24.73% |
| Major | 96 | 4 | 42 | 30 | 20 | 45.65% | 32.61% | 21.74% |
| LEAM | 95 | 7 | 59 | 16 | 13 | 67.05% | 18.18% | 14.77% |
| BERT | 88 | 8 | 37 | 24 | 19 | 46.25% | 30.00% | 23.75% |
| BugFarm (GPT-4o) | 94 | 9 | 62 | 11 | 12 | 72.94% | 12.94% | 14.12% |
| LLMopheus (GPT-4o) | 95 | 11 | 26 | 44 | 14 | 30.95% | 52.38% | 16.67% |
| LLMut (GPT-3.5) | 95 | 12 | 66 | 12 | 5 | 79.52% | 14.46% | 6.02% |
| LLMut (GPT-4o) | 95 | 11 | 53 | 16 | 15 | 63.10% | 19.05% | 17.86% |
| LLMut (GPT-4o-M) | 95 | 12 | 51 | 15 | 17 | 61.45% | 18.07% | 20.48% |
| LLMut (DS-761b) | 95 | 12 | 45 | 16 | 22 | 54.22% | 19.28% | 26.51% |
| LLMut (DS-236b) | 95 | 13 | 54 | 15 | 13 | 65.85% | 18.29% | 15.85% |
| LLMut (SC-16b) | 93 | 16 | 62 | 11 | 4 | 80.52% | 14.29% | 5.19% |
| LLMut (CL-13b) | 96 | 11 | 68 | 12 | 5 | 80.00% | 14.12% | 5.88% |
As shown in Table 13, the distributions of alive mutations differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutations, respectively. First, LLM-based methods exhibit a substantially higher proportion of (i.e., not covered) mutations, meaning that more of their alive mutations reside in previously uncovered code lines. For instance, among the alive mutations generated by GPT-3.5 and BugFarm, 79.52% and 72.94% fall into the category, respectively, whereas PIT and Major exhibit much lower proportions of 46.24% and 45.65%. This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas. Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutations compared to LLM-based methods. For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas all LLMs based on our prompt remain below 20%. These types of alive mutations require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage. Therefore, the overall implication is that LLM-generated alive mutations are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.
4. Discussion
In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
4.1. Sensitivity to Experimental Settings
We discuss the experiment setups that may affect the evaluation results, particularly the setups related to LLMs.
4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutation generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the recent open-source model DeepSeek-V3-761b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get five types of prompts, shown as follows:
Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of the corresponding unit tests of the target method into Context.
Prompt 5 (P5): P1 with Whole Java File. P5 is a variant of P1, which uses the code of the entire Java file as the code context, rather than the entire method.
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||||||
| P1 | P2 | P3 | P4 | P5 | P1 | P2 | P3 | P4 | P5 | |
| Mut. # | 3681 | 3688 | 3350 | 3210 | 3496 | 3851 | 3847 | 3834 | 3844 | 3878 |
| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.694 | 0.765 | 0.750 | 0.760 | 0.753 | 0.776 |
| Real Bug Detec. | 79.5% | 78.4% | 72.2% | 72.5% | 79.5% | 85.5% | 85.8% | 83.9% | 84.7% | 83.1% |
| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 62.3% | 67.5% | 65.6% | 66.4% | 65.8% | 66.4% |
| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 47.3% | 53.2% | 51.7% | 50.9% | 50.0% | 49.1% |
| Compling Rate | 66.9% | 66.8% | 60.0% | 61.2% | 70.3% | 82.8% | 83.5% | 78.0% | 82.0% | 79.5% |
| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 16.9% | 7.3% | 6.4% | 8.5% | 7.0% | 8.2% |
| Avg. Gen. Time | 1.789 | 1.782 | 1.09 | 1.884 | 1.907 | 2.573 | 2.571 | 2.488 | 2.601 | 2.596 |
| Avg. # Token | 105 | 95 | 44 | 219 | 393 | 104 | 92 | 43 | 219 | 392 |
Note: The prompt template that achieves the best performance is highlighted in grey background color.
Table 14 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek-V3-761b across multiple metrics. Across both models, P1 delivers the best overall performance on all effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), demonstrating its strength in generating mutations that are semantically closer to real bugs. For the validity metrics, P1 ranks second in compilation rate on GPT-3.5 and first on DeepSeek-V3-761b, while achieving the lowest duplicate mutation rate on GPT-3.5 and third-lowest on DeepSeek-V3-761b. Notably, compared with P5, which provides the entire file as context, P1 relies only on the full method yet achieves higher semantic similarity to real bugs, with reduced generation time and token usage, while maintaining comparable compilation and duplicate mutation rates.
Across P1 to P5, we observe a clear context–performance trade-off. Removing relevant context (P3) substantially degrades effectiveness, as seen in lower real bug detectability, average Ochiai, and coupling rates. Providing a focused, method-level context (P1 and P2) delivers the best overall balance of effectiveness and validity. However, adding excessive context beyond the focal method (P4 and P5) does not yield further improvements and often increases token usage and generation time, while in some cases even reducing key effectiveness metrics. This suggests that moderate, relevant context benefits LLM mutation generation, whereas both insufficient and excessive context hurt efficiency and quality.
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 16. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05. Although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
4.1.2. Experiment Setup of the Context Length
| Metric | 1-Line Context | 2-Line Context | 3-Line Context |
| Real Bug Dectc. | 85.0% | 83.3% | 83.3% |
| Avg. Ochiai | 33.3% | 35.0% | 30.0% |
| Coupling Rate | 36.6% | 35.2% | 37.1% |
| Compilability Rate | 62.9% | 63.2% | 67.2% |
| Dup. Mut. Rate | 10.8% | 12.5% | 11.9% |
| Settings | Spea. Coef | -value | Pear. Coef | -value |
| GPT-3.5 P1 v.s. GPT-3.5 P2 | 0.9762 | 3.31E-05 | 0.9999 | 1.35E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P3 | 0.9040 | 2.00E-03 | 0.9999 | 8.66E-23 |
| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.94E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P5 | 0.9762 | 3.31E-05 | 0.9999 | 1.57E-26 |
| DS-671b P1 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
| DS-671b P1 v.s. DS-671b P3 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
| DS-671b P1 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
| DS-671b P1 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
| 1-Line vs. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
| 2-Line vs. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
| 1-Line vs. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
| QB-6a (Default) v.s. QB-6b | 0.9429 | 4.81E-03 | 0.9984 | 1.00E-08 |
| QB-6a (Default) v.s. QB-3 | 0.9856 | 3.09E-05 | 0.9999 | 1.62E-13 |
| QB-6a (Default) v.s. QB-4 | 0.9643 | 4.54E-04 | 0.9999 | 7.28E-23 |
| QB-6a (Default) v.s. QB-5 | 0.9643 | 4.54E-04 | 0.9999 | 9.62E-24 |
| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 9.39E-16 |
| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 3.34E-15 |
| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 5.01E-14 |
| QB-6a (Default) v.s. QB-10 | 0.9856 | 3.09E-04 | 0.9999 | 2.89E-13 |
| QB-6a (Default) v.s D4J-6 | 0.9524 | 2.60E-04 | 0.9980 | 2.13E-08 |
QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 15. Similarly, the middle part of Table 16 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95, and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
4.1.3. Impact of Different Few-Shot Examples
| Metric | QB-6a | QB-3 | QB-4 | QB-5 | QB-6b | QB-7 | QB-8 | QB-9 | QB-10 | D4J-6 |
| Real Bug Detec | 0.917 | 0.933 | 0.883 | 0.883 | 0.917 | 0.900 | 0.900 | 0.917 | 0.867 | 0.900 |
| Avg. Ochiai | 0.467 | 0.450 | 0.400 | 0.433 | 0.433 | 0.467 | 0.433 | 0.433 | 0.400 | 0.367 |
| Coupling Rate | 0.136 | 0.128 | 0.126 | 0.127 | 0.132 | 0.112 | 0.113 | 0.121 | 0.136 | 0.139 |
| Comp. Rate | 67.5% | 62.7% | 65.6% | 66.8% | 63.9% | 65.1% | 66.0% | 61.8% | 64.9% | 62.1% |
| Dup.Mut.Rate | 13.7% | 12.8% | 12.6% | 9.4% | 12.8% | 11.4% | 13.2% | 13.8% | 13.6% | 13.9% |
| Avg. # Token | 106 | 98 | 102 | 103 | 106 | 108 | 109 | 110 | 113 | 107 |
QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 17.
We further compute their similarity, shown as the bottom part of Table 16. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the -values are all below 0.05. Therefore, despite variations in few-shot examples, there is no statistically significant evidence indicating superior performance among the prompts.
4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397. This ensures a fair comparison across approaches by controlling for differences in mutation counts, thereby eliminating the potential bias that methods generating more mutations might appear to perform better simply due to quantity rather than quality. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 18. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
| Name | R. B. Detec. | Ochiai Coeffi. | Coup. Rate | Comp. Rate | Dup. Rate |
| PIT | 31.6% | 12.6% | 17.4% | — | — |
| Major | 68.8% | 22.2% | 45.4% | 96.9% | 0.0% |
| LEAM | 64.6% | 14.1% | 37.2% | 54.1% | 2.6% |
| BERT | 66.0% | 19.0% | 43.6% | 26.1 | 1.9% |
| BugFarm (GPT-4o) | 58.6% | 17.7% | 47.9% | 74.2% | 7.5% |
| LLMorpheus (GPT-4o) | 67.5% | 23.6% | 50.2% | 68.2% | 7.3% |
| LLMut (GPT-3.5) | 82.2% | 26.5% | 45.6% | 61.2% | 9.9% |
| LLMut (GPT-4o) | 87.4% | 36.0% | 48.8% | 75.5% | 7.7% |
| LLMut (GPT-4o-M) | 88.6% | 31.8% | 47.9% | 74.3% | 7.1% |
| LLMut (DS-761b) | 89.2% | 39.4% | 49.4% | 76.3% | 7.7% |
| LLMut (DS-236b) | 87.0% | 34.1% | 47.9% | 74.6% | 7.6% |
| LLMut (SC-16b) | 33.9% | 15.4% | 27.1% | 12.4% | 9.6% |
| LLMut (CL-13b) | 68.5% | 16.9% | 43.6% | 69.5% | 38.3% |
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||
| Default | More Detailed | Default + 10 Err. | Default | More Detailed | Default + 10 Err. | |
| Real Bug Detec | 0.917 | 0.900 | 0.900 | 0.950 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.500 | 0.500 | 0.517 |
| Coupling Rate | 0.136 | 0.129 | 0.117 | 16.9% | 17.8% | 16.0% |
| Comp. Rate | 67.5% | 71.1% | 68.7% | 78.4% | 76.4% | 77.5% |
| Dup.Mut.Rate | 13.7% | 14.1% | 10.4% | 5.1% | 4.1% | 6.2% |
Note: The prompt achieving the best performance in terms of a metric is highlighted in grey background color.
4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
In software engineering practice, even when using the same information sources, users may design prompts differently, for example, varying the level of detail in the requirements description or introducing minor language errors. To investigate the impact of such variations, we derived two prompt variants from the default prompt shown in Figure 1.
More Detailed Output Specification: replacing the output specifications of the default prompt by explicitly requiring: Prohibit generating identical mutants, i.e., ensure ‘aftercode’ ‘precode’, and all mutants are unique. Write all output mutants in a JSON file.
Default + 10 Errors: randomly injecting 10 English grammar mistakes into the default prompt, such as subject–verb agreement errors (e.g., it are) and noun number errors (e.g., two mutation).
Following the same experimental setup as before, we compared these two variants using GPT-3.5 and DeepSeek-V3-761b, with the results summarized in Table 19.
For GPT-3.5, the More Detailed Output Specification variant slightly improves the compilation rate (+3.6%) but also increases the duplicate mutation rate. In contrast, the Default + 10 Errors variant reduces real bug detection and coupling rate compared to the default prompt. Interestingly, the Default + 10 Errors variant, presumably the lowest in prompt quality, achieves the best performance on duplicate mutation rate. For DeepSeek-V3-761b, the differences across variants are minimal. For instance, the default and the More Detailed Output Specification prompts achieve identical real bug detection and average Ochiai, and the More Detailed Output Specification variant even shows a small increase in average Ochiai.
Overall, these results suggest that prompt wording has a certain degree of randomness in its effect on mutation generation performance. To further verify this, we conducted similarity analysis across the three prompts for each metric and found no statistically significant superiority of any one prompt over the other two.
| Metric | LLMut (GPT-3.5) | LLMut (GPT-4o) | LLMut (DS-671b) | |||
| FOM | HOM | FOM | HOM | FOM | HOM | |
| Real Bug Detec. | 0.917 | 0.900 | 0.933 | 0.933 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.483 | 0.500 | 0.483 |
| Coupling Rate | 0.136 | 0.249 | 0.151 | 0.307 | 0.169 | 0.306 |
| Compilability Rate | 67.5% | 63.4% | 76.3% | 70.2% | 78.4% | 73.3% |
| Dup. Mut. Rate | 13.7% | 5.2% | 5.3% | 3.5% | 5.1% | 3.1% |
Note: The order achieving the best performance in terms of a metric is highlighted in grey background color.
4.3. Effectiveness on High-Order Mutation Generation
Previously, we only evaluated LLMs in generating first-order mutations (FOM). To explore the effectiveness of LLMs in generating high-order mutations (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating the recent mainstream models, including GPT-3.5, GPT-4o, and DeepSeek-V3-761b.
We modified the original prompt in Figure 1 to explicitly request HOM generation. Table 20 shows the results. For the effectiveness metrics, HOMs generally achieve average Ochiai scores and higher coupling rates across these models, indicating closer semantic alignment with real bugs. However, FOMs achieve higher real bug detectability on all models. For the validity metrics, HOMs exhibit lower duplicate mutation rates, suggesting improved mutation diversity. However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors. Our preliminary experiments suggest that strengthening LLM-based HOM generation is promising, as HOMs are semantically closer to real bugs and face more challenges in maintaining syntactic correctness.
4.4. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches and suggest potential improvements.
4.4.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations. For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations. In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
Additionally, our study explores how different prompt designs, prompt turbulence, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
Moreover, our study demonstrates that LLMs have clear potential in generating HOMs. Compared with first-order mutations (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability. The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
4.4.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
4.4.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
Moreover, exploring the applications of LLM-generated mutation in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL). Our study has shown that LLM-generated mutations often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
Another important direction is cross-language generalization. While our study primarily focuses on Java, mutation testing is equally critical in languages such as C/C++, Python, JavaScript, and Rust. Traditional mutation testing tools usually require language-specific implementations, leading to duplicated engineering efforts and limited portability. In contrast, modern LLMs possess inherent multilingual capabilities, as they are pre-trained on large-scale, cross-language code corpora. This enables them to naturally transfer mutation strategies across programming languages without retraining from scratch. Future work should investigate how well LLMs can exploit this cross-language knowledge, and how prompts, training data, or filtering strategies can be adapted to different language paradigms (e.g., strongly typed vs. dynamically typed). Advancing towards multilingual mutation generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
4.5. Limitations of Our Study
Our study has the following limitations that are left for further exploration.
First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators. For instance, the phenomenon of subsuming mutations (Garg et al., 2022; Diniz et al., 2021), where one mutation is always killed by any test that kills another, has been studied in the mutation testing literature as a way to assess better redundancy and the true fault-detection power of test suites. Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutations beyond the metrics reported in this paper. Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
Second, our experiments are limited to the Java programming language. Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms. Mutation behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments. Future research should examine the generalizability of LLM-based mutation generation across multiple languages. Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutation generation a promising direction for broadening the applicability of our findings.
Third, our study relies on fixed few-shot examples, which restricts LLMs to producing context-specific mutations for the given code snippets. In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators. Consequently, it is still an open question how LLMs would perform if constrained to the same operator set as rule-based approaches.
5. Threats to Validity
In this section, we illustrate the threats to the validity of our study.
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek). Moreover, our study further adopts two state-of-the-art LLM-based mutation generation approaches (i.e., BugFarm and LLMorpheus) as baselines.
The metrics we used may threaten the conclusion of our study. Each metric has inherent limitations. For instance, Real Bug Detection Rate and Coupling Rate capture only partial aspects of similarity to real bugs: a method may detect many bugs but with weak coupling on each bug, or achieve strong coupling concentrated on a few bugs. Likewise, AST Edit Distance reflects structural differences, but a higher syntactic distance does not necessarily imply larger semantic changes. Additionally, the measurement of Average Generation Time may be affected by external factors such as GPU type, service provider implementation, and network fluctuations, introducing unavoidable noise. To mitigate the bias of individual metrics, we report a comprehensive set of complementary measures. Particularly, besides time cost, we also report token consumption, which provides a more stable and model-independent estimate of computational and financial cost.
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of the most LLMs we use (expect DeepSeek-V3-761b only) and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek-V3-761b (72), DeepSeek-V2-236b (66), BugFarm (59), LLMorpheus (26), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek-V3-761b (35), DeepSeek-V2-236b (32), BugFarm (22), LLMorpheus (19), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompt design, prompt turbulence, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
The subjective nature of human decisions in labeling for equivalent mutations, non-compilation errors, and alive mutations is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
6. Related Work
In this section, we introduce the related work of our study.
6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to be coupled with real faults/bugs to allow these techniques to be effective.
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. LEAM++, the extension of LEAM, further explores learning-based mutation selection (Tian et al., [n. d.]). However, these approaches suffer from the challenges in generating syntactically correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022). BERT generates mutation by masking the code elements to be mutated and then predicting the mutation via BERT (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently investigated the use of LLMs for mutation generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutations. In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches. Furthermore, we explore different prompt strategies, such as incorporating few-shot examples and enforcing structured JSON output, and we systematically examine multiple prompt templates in the context of mutation testing. Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al. (Dakhel et al., 2024) and Ibrahimzada et al. (Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. BugFarm (Ibrahimzada et al., 2023) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.
6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023b). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo, GPT-4o, and GPT-4o-Mini are closed-source LLMs released by OpenAI (Achiam et al., 2023), while DeepSeek-Coder, StarChat, and CodeLlama are open-source. An open-source LLM is usually fine-tuned against a base model for specified tasks. DeepSeek-Coder (Zhu et al., 2024) is based on DeepSeek. CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets. StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Particularly, we evaluate not existing LLM-based approaches but also our new prompt for mutation generation. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.